GNU Interactive Tools

A Set of Interactive Programs Edition 2.9.4, for GNUIT version 4.9.5

January 2008

by Tudor Hulubei, Andrei Pitis and Ian Beckwith

GNUIT: A set of interactive tools, by Tudor Hulubei and Andrei Pitis. This file documents the GNU Interactive Tools package. Copyright (C) 1993-1998, 2006-2008 Free Software Foundation, Inc.

Permission is granted to copy, distribute and/or modify this document under the terms of the GNU Free Documentation License, Version 1.3 or any later version published by the Free Software Foundation; with no Invariant Sections, with no Front-Cover Texts, and with no Back-Cover Texts. A copy of the license is included in the section entitled “Copying This Manual”.

1 Introduction

GNUIT is a set of interactive tools. It contains an extensible file system browser, an ascii/hex file viewer, a process viewer/killer and some other related utilities and shell scripts. It can be used to increase the speed and efficiency of most of the daily tasks such as copying and moving files and directories, invoking editors, compressing and uncompressing files, creating and expanding archives, compiling programs, sending mail, etc. It looks nice, has colors (if the standard ANSI color sequences are supported) and is user-friendly.

GNUIT runs on a wide variety of UNIX systems because it uses the GNU Autoconf package to get system specific information. Please refer to the PLATFORMS file included in the standard distribution for a detailed list of systems on which GNUIT has been tested.

One of the main advantages of GNUIT is its flexibility. It is not limited to a given set of commands. The configuration file can be easily enhanced, allowing the user to add new commands or file operations, depending on its needs or preferences.

GNUIT also provides a shell like command prompt, just to make sure that the entire power of the UNIX shell commands is still there.

2 Distributing GNU Interactive Tools

GNUIT is "free software"; this means that everyone is free to use it and free to redistribute it on certain conditions. GNUIT is not in the public domain; it is copyrighted and there are restrictions on its distribution, but these restrictions are designed to permit everything that a good cooperating citizen would want to do. What is not allowed is to try to prevent others from further sharing any version of GNUIT that they might get from you. The precise conditions are found in the GNU General Public License that comes with GNUIT and also appears following this section.

The easiest way to get a copy of GNUIT is from someone else who has it. You need not ask for our permission to do so, or tell any one else; just copy it. If you have access to the Internet, you can get the latest distribution version of GNUIT from host ‘ftp.gnu.org’ using anonymous login. See the file ‘/pub/gnu/GETTING.GNU.SOFTWARE’ on that host to find out about your options for copying and which files to use.

You may also receive GNU Interactive Tools when you buy a computer. Computer manufacturers are free to distribute copies on the same terms that apply to everyone else. These terms require them to give you the full sources, including whatever changes they may have made, and to permit you to redistribute the GNU Interactive Tools received from them under the usual terms of the General Public License. In other words, the program must be free for you when you get it, not just free for the manufacturer.

3 Using GNU Interactive Tools

The GNUIT package contains three interactive programs and a few additional utilities. Here there is a description of each of them.

3.1 The GIT file system browser

gitfm is a file system browser with some shell like features designed to make your work much easier and much efficient. It displays one or two panels, each one containing a file system directory. You can browse the directory tree with the usual cursor keys, pressing ENTER when you want to enter or leave a directory and TAB when you want to change the panels.

Under the two panels there is a shell like input line which you can use to type normal shell commands. The input line can handle an unlimited number of characters and keeps a history of typed commands (using the GNU history library).

Under the input line there is a status bar. You can see there the status of the currently executed command, the warnings and errors and you will be prompted if a decision has to be taken.

3.1.1 Key binding conventions

gitfm now follows a new, easy to remember, scheme to bind commands on keys. This is only a convention, if you define new key bindings you may, or may not follow it.

All the file commands start with ^C. This prefix can be followed by some modifiers, in order to affect the default behavior of the given command. These modifiers are b and r.

b - this modifier specifies that the command will run in background:

     ^CM = CHMOD; chmod %s{New mode of %i: ,%m} %i;;;;y

defines a command that changes the current selected files mode in foreground, while

     ^CbM = B-CHMOD; chmod %s{New mode of %i: ,%m} %i&;;;;y

defines a background command that does the same thing.

r - this modifier specifies that the command will be run recursively:

     ^CrM = R-CHMOD; chmod -R %s{New mode of %i: ,} %i;;;;y

defines a command that recursively changes the mode of the selected entries.

The b and r modifiers can be combined, the resulting command running recursively and in background:

     ^CbrM = B-R-CHMOD; chmod -R %s{New mode of %i: ,} %i&;;;;y

You should also note that for some commands (like gzip) there is no need for a nonrecursive version. Running gzip recursively on files is harmless. If there is a directory between these files, gzip will recursively compress that directory, so you can use the same key binding for recursively and non-recursively compressing. In fact, it is a matter of selecting files or directories.

Unfortunately, we can’t run chmod recursively trying to change the mode of all the files in a directory to 0644 because that directory might contain subdirectories and removing the execution permission from them is a bad idea. So, in this case, we need separate commands.

3.1.2 Command line

This is a brief description of the command line arguments. -h print this help message -v print the version number -c use ANSI colors -b don’t use ANSI colors -l don’t use the last screen character -p output final path at exit

The -p option can be used to make gitfm force bash (assuming that you’re using it as your shell) chdir to the last directory gitfm was in before quitting. In order to do this, you need to invoke gitfm using this function (put it into your .profile):

function g

     {

gitfm -p $ 3> /tmp/gitfm.p.$$

if test -s /tmp/gitfm.p.$$; then if test -d "‘cat /tmp/gitfm.p.$$‘"; then

cd "‘cat /tmp/gitfm.p.$$‘" else cd

fi fi

rm -f /tmp/gitfm.p.$$

     }

This will not work if you suspend gitfm. Nothing bad will happen, just the chdir will not be performed.

3.1.3 Panel modes

gitfm has three major modes of displaying the panels. In the first (default) mode, two panels are displayed, each one using half of the screen. In the second mode, only one panel uses the entire screen. In the third mode, only the status bar and the input line are displayed, both panels being hidden.

Briefly, a panel can use the entire screen or just half of it. Even when a panel is hidden, it still exists.

Users can switch between these three major modes as needed:

^X 0

Enlarges the other panel to use the entire screen. It also changes the minor mode to ‘Enable all’. The current panel will become

     invisible (‘enlarge-other-panel’).
^X 1

Enlarges the current panel to use the entire screen. It also changes the minor mode to ‘Enable all’. The other panel will become

invisible (‘enlarge-panel’).

^X 2

Switches back to the two panel mode (‘two-panel-mode’).

^O, ESC o

Switches to the tty mode (no panels on the screen) (‘tty-mode’). A panel displays the files and subdirectories in a directory. You can optionally specify

some additional information about each entry (file, directory, …) to be displayed (a minor mode).

When using the full screen mode, all the minor modes here can be used. In half screen mode, the ‘panel-enable-all’ mode is not available.

These are the panel minor modes:

ESC e o

Display the entry owner and group (‘panel-enable-owner-group’).

ESC e d

Display the entry date and time (‘panel-enable-date-time’).

ESC e s

Display the entry size (‘panel-enable-size’).

ESC e S

Display the entry size, scaled (e.g. ‘123M’) (‘panel-enable-abbrevsize’).

ESC e m

Display the entry mode (‘panel-enable-mode’).

ESC e f

Display the entry full name (‘panel-enable-full-name’).

ESC e a

Display the entire information about file (‘panel-enable-all’). This mode is only available if the panel has been enlarged to use the

     entire screen with ‘enlarge-panel’ or ‘enlarge-other-panel’
     (‘panel-enable-all’).

There is another way of changing the panel minor modes:

^], ^[]

Switches to the next panel minor mode (‘panel-enable-next-mode’).

3.1.4 Sorting methods

Entries in a panel can be sorted in different ways. These are the available options:

ESC s n

Display the panel entries sorted by their names

     (‘panel-sort-by-name’).
ESC s e

Display the panel entries sorted by their extensions

     (‘panel-sort-by-extension’).
ESC s s

Display the panel entries sorted by their sizes

(‘panel-sort-by-size’).
ESC s d

Display the panel entries sorted by their ‘last modified’ stamps

(‘panel-sort-by-date’).
ESC s m

Display the panel entries sorted by their modes

(‘panel-sort-by-mode’).
ESC s o i

Display the panel entries sorted by their owner ids

(‘panel-sort-by-owner-id’).
ESC s g i

Display the panel entries sorted by their group ids

(‘panel-sort-by-group-id’).
ESC s o n

Display the panel entries sorted by their owner names

(‘panel-sort-by-owner-name’).
ESC s g n

Display the panel entries sorted by their group names

(‘panel-sort-by-group-name’).

There is also another way to change the sort method:

ESC s u

Switch to the next panel sort method (‘panel-sort-next-method’).

3.1.5 Moving the cursor in the panel

Moving the cursor in the panel is very easy. If your keyboard has arrows, use them. If the arrow keys don’t work (it might be due to a badly configured TERM environment variable), you can use the Emacs commands bindings as well.

UP, ^P

Move the cursor vertically up one entry (‘previous-line’).

DOWN, ^N

Move the cursor vertically down one entry (‘next-line’).

HOME, ESC <

Move the cursor on the first entry in the panel

(‘beginning-of-panel’).
END, ESC >

Move the cursor on the last entry in the panel (‘end-of-panel’).

PGUP, ESC v

Move the cursor vertically down one page (‘scroll-down’).

PGDOWN, ^V

Move the cursor vertically down one page (‘scroll-up’).

ESC g

Scroll the panel entries to the left (‘horizontal-scroll-left’).

ESC j

Scroll the panel entries to the right (‘horizontal-scroll-right’).

^X P

In order to optimize the screen output, you can modify the scroll step (‘set-scroll-step’). This is the number of lines to try scrolling a panel when the cursor moves out. The ‘StartupScrollStep’ specifies the initial scroll step, but using ‘set-scroll-step’ you can dynamically change it.

TAB, ^I, ^X o

Move the cursor in the other panel (‘other-panel’).

^X P

Switch the two panels. This command works even when gitfm is not

in the ‘two panels’ mode (‘switch-panels’).

3.1.6 Selecting files

INS, ^T, ^X \, ^\

Toggle the ‘selected’ flag of the current entry

(‘select-entry’).
^C s

Select (marks) all the files matching at least one pattern from a space separated list of shell patterns. Spaces and \s are allowed in the patterns but they have to be escaped with a \. The user will be prompted for a pattern to match against

(‘select-files-matching-pattern’).
^C u

Unselect (unmarks) all the files matching at least one pattern from a space separated list of shell patterns. Spaces and \s are allowed in the patterns but they have to be escaped with a \. The user will be prompted for a pattern to match against

(‘unselect-files-matching-pattern’).
ESC +

Select (marks) all the files having the same extension as the current file. If the current file name doesn’t have an extension or starts with a dot, no files are selected (‘select-extension’).

ESC -

Unselect (unmarks) all the files having the same extension as the current file. If the current file name doesn’t have an extension or starts with a dot, no files are unselected (‘unselect-extension’).

See Section 3.1.8.8 [Selecting Files], page 10, for additional ways of selecting and unselecting files.

3.1.7 Incremental searching files in a panel

Users sometime need to search a file in a panel, especially when the panel contains a big number of entries. For that reason gitfm provides an incremental search feature. Using forward and backward incremental search, files can be very easy located. Wrapped incremental search is also provided.

^S, ^Xs

Incremental search forward a file name in the current panel (‘isearch-forward’). Pressing ^S or ^Xs again will force gitfm to go to the next entry that matches the current isearched string. When the end of the panel is reached, the isearch is restarted from its beginning.

^R, ^Xr

Incremental search backward a file name in the current panel (‘isearch-backward’). Pressing ^R or ^Xr again will force gitfm to go to the next entry that matches the current isearched string. When the beginning of the panel is reached, the isearch is restarted from its end.

3.1.8 Using the input line

The input line is one of the main methods used by gitfm to interact with the user. All the answers the user should give in order to perform some operation and all the shell like commands are built using it. So here is a description of all the basic editing operations that the ‘input line’ provides. They are very much inspired from Emacs, so Emacs users should have no problem using them.

3.1.8.1 Inserting Text

Typing characters is the most usual way of inserting text into the input line. Key sequences starting with printable ascii characters are not allowed in gitfm so typing a for example results in inserting a at the current point position. Of course, there are some other ways of inserting text into the command line and here there is a description of most of them.

ESC RET

Copy the current entry name into the input line at the current point

     position (‘entry-to-input-line’).
ESC ESC RET

Copy the other panel path into the input line at the current point

     position (‘other-path-to-input-line’).
^X ^I

Copy the names of all the selected entries into the input line at the

     current point position (‘selected-entries-to-input-line’).

3.1.8.2 Moving Point

^B, LEFT

Move the point backward one character (‘backward-char’).

^F, RIGHT

Move the point forward one character (‘forward-char’).

ESC b

Move the point one word backward (‘backward-word’).

ESC f

Move the point one word forward (‘forward-word’).

^A

Move the cursor at the beginning of the input line (‘beginning-of-line’).

^E

Move the cursor at the end of the input line (‘end-of-line’).

3.1.8.3 Deleting and killing text

DEL, ^D

Delete the character under the cursor (‘delete-char’).

^H, BKSPC

Delete the character before the cursor (‘backward-delete-char’).

ESC BKSPC

Delete backward one word (‘backward-kill-word’).

ESC d

Delete forward one word (‘kill-word’).

ESC k

Delete the entire line (‘kill-line’).

^U

Delete all the characters between the beginning of the input line and

the point (‘kill-to-beginning-of-line’).
^K

Delete all the characters between the point and the end of the input

line (‘kill-to-end-of-line’).
ESC SPC

Delete all the spaces around the point, leaving only one space

(‘just-one-space’).
ESC \

Delete all the spaces around the point

(‘delete-horizontal-space’).
^W

Save the region between the point and the mark into the kill "ring" and then kills it (‘kill-region’). Note that there is no real kill-ring here. The so-called kill-ring has only one entry.

ESC w

Save the region between the point and the mark without killing it

(‘kill-ring-save’).

3.1.8.4 Case conversion of words.

ESC l

Convert the following word to lower case, moving over.

     (‘downcase-word’).
ESC u

Convert the following word to upper case, moving over.

     (‘upcase-word’).
ESC c

Capitalize the following word, moving over.

     (‘capitalize-word’).

3.1.8.5 Reusing recent input line arguments

A separate history is kept for both built-in and user-defined commands. If you call a command that you have used before, you can re-edit a previously entered string in order to minimize the amount of characters needed to be typed for the new one. There is no limit on the number of strings that can be kept in the history.

ESC p

Walk backward through the history of previously entered strings

     (‘previous-history-element’).
ESC n

Walk forward through the history of previously entered strings

     (‘next-history-element’).

3.1.8.6 Commands to set the mark

^SPC

Set the mark at the current point position (‘set-mark’).

^X ^X

Exchange the current point position with the mark one

     (‘exchange-point-and-mark’).

3.1.8.7 Reinserting recently killed text

^Y

Reinsert a previously killed text at the current point position (‘yank’).

3.1.8.8 Selecting files matching patterns

If the very first character in the input line is a ‘+’, what comes after it is considered a (space separated) list of shell patterns, and all the files that match at least one pattern from that list will be marked as selected. An empty list of shell patterns (i.e. the ‘+’ by itself) will cause all the files to be selected.

If the very first character in the input line is a ‘-’, the space separated list of shell patterns that follows is used to unselect files. An empty list of shell patterns (i.e. the ‘-’ by itself) will cause all the selected files to be unselected.

Finally, if the first and only character in the input line is a ‘*’, then all the selected files will become unselected, and all the unselected files will become selected.

3.1.9 File operations

3.1.9.1 Copying Files

F5, ESC 5, ^C C

Copy the currently selected entries to the user supplied path (‘copy’).

^C b C

Copy the currently selected entries to the user supplied path. The operation is performed in background (‘B-COPY’).

3.1.9.2 Moving Files

F6, ESC 6, ^C T

Move the currently selected entries to the user supplied path (‘move’).

^C b T

Move the currently selected entries to the user supplied path. The operation is performed in background (‘B-MOVE’).

3.1.9.3 Creating Files

The easiest way to create a new file is to start an editor passing the file name as an argument. Most editors will try to create the file if the file doesn’t exist. See Section 3.1.9.10 [Editing Files], page 14, for more information.

3.1.9.4 Deleting Files

F8, ESC 8, ^C D

Delete the currently selected entries (‘delete’).

^C b D

Delete the currently selected entries. The operation is performed in background (‘B-DELETE’).

3.1.9.5 Linking Files

^C H

Create a hard link from the current files to a user supplied file name (‘LINK’).

^C b H

Create a hard link from the current files to a user supplied file name (‘B-LINK’). The action is performed in background.

^C S

Create a symbolic link from the current files to a user supplied file name (‘SYMLINK’).

^C b S

Create a symbolic link from the current files to a user supplied file name (‘B-SYMLINK’). The action is performed in background.

3.1.9.6 Renaming Files

^C R

Rename the current file or directory with the user supplied name

(‘RENAME’).
^C b R

Rename the current file or directory with the user supplied name. The operation is performed in background (‘B-RENAME’).

^C n d

Change the name of all the selected entries to lowercase.

(‘name-downcase’).
^C n u

Change the name of all the selected entries to uppercase.

(‘name-upcase’).

3.1.9.7 Splitting files into smaller parts

^C /

Split the current file into several smaller files of a given size and named based on a given prefix (‘SPLIT’).

^C b /

Split the current file into several smaller files of a given size and named based on a given prefix. The operation is performed in background

(‘B-SPLIT’).

3.1.9.8 Packing files into the minimum number of bins

^C ~

Pack the files into the smallest number of bins. This is an approximation - the problem is NP-complete and no known algorithm can guarantee a solution better than ‘(11/9) * OPTIMAL + 4’.

To make things even worse, for large files, there is no portable way to predict how many blocks the file system implementation will require for indirect blocks, directories, etc. So keep in mind that this is only an approximation.

Bin packing can be useful when you want to put a bunch of files on floppies or zip disks and you want to optimize things a little bit

(‘bin-packing’).

‘gitfm’ assumes that you want to pack all the files in the current directory - if there is any selected file in that directory it will be unselected first. Then ‘gitfm’ will ask for a bin size, and select the files that should go in the first bin. You are supposed to place those files in the first bin (e.g. a tar archive), remove them from the current directory, then run ‘bin-packing’ again, to obtain the list of the files that should go into the second bin, etc.

3.1.9.9 Changing a file’s mode, owner and group

^C M

Change the mode of the currently selected entries (‘CHMOD’).

^C b M

Change the mode of the currently selected entries. The operation is performed in background (‘B-CHMOD’).

^C r M

Recursively change the modes of the selected entries if one of them is a directory (‘R-CHMOD’).

^C b r M

Recursively change the modes of the selected entries if one of them is a directory. The operation is performed in background (‘B-R-CHMOD’).

^C O

Change the owner of the currently selected entries (‘CHOWN’).

^C b O

Change the owner of the currently selected entries. The operation is performed in background (‘B-CHOWN’).

^C r O

Recursively change the owners of the selected entries if one of them is a directory (‘R-CHOWN’).

^C b r O

Recursively change the owners of the selected entries if one of them is a directory. The operation is performed in background

(‘B-R-CHOWN’).
^C G

Change the group of the currently selected entries (‘CHGRP’).

^C b G

Change the group of the currently selected entries. The operation is performed in background (‘B-CHGRP’).

^C r G

Recursively change the groups of the selected entries if one of them is a directory (‘R-CHGRP’).

^C b r G

Recursively change the groups of the selected entries if one of them is a directory. The operation is performed in background

(‘B-R-CHGRP’).

3.1.9.10 Editing Files

F4, ESC 4

Call the default editor with the current file name as an argument (‘EDIT’).

^X e

Call the default editor with the selected entry names as arguments

(‘MULTIPLE-EDIT’).
^X ^F

Create a new file by calling the default editor with the user supplied file name as an argument (‘FILE-CREATE’).

^X 4 a

Call the default editor in order to edit the ‘ChangeLog’ file

(‘CHANGE-LOG’).

The default editor can be specified using the EDITOR or GNUIT EDITOR environment variables. See Section 4.1 [Environment Variables], page 33, for more information.

$GNUIT EDITOR used to be called $GIT EDITOR. The old name is still accepted for backwards-compatibility.

3.1.9.11 Viewing Files

F3, ESC 3

Call the default viewer (gitview) with the current file name as argument (‘VIEW’).

^X v

Call the default pager (more) with the currently selected entry names as arguments (‘MULTIPLE-VIEW’).

3.1.9.12 Compressing Files

^C z

Compress the currently selected entries with gzip

(‘COMPRESS’).
^C b z

Compress the currently selected entries with gzip. The operation is performed in background (‘B-COMPRESS’).

^C Z

Uncompress the currently selected entries with gunzip

(‘UNCOMPRESS’).
^C b Z

Uncompress the currently selected entries with gunzip. The operation is performed in background (‘B-UNCOMPRESS’).

^C f Z

Uncompress the currently selected entries with gunzip (‘F-UNCOMPRESS’). Force uncompression of links.

^C b f Z

Uncompress the currently selected entries with gunzip. Force uncompression of links. The operation is performed in background

(‘BF-UNCOMPRESS’).
^C .

Compress the currently selected entries with bzip2

(‘BZIP2-COMPRESS’).
^C b .

Compress the currently selected entries with bzip2. The operation is performed in background (‘B-BZIP2-COMPRESS’).

^C o

Uncompress the currently selected entries with bunzip2

(‘BZIP2-UNCOMPRESS’).
^C b o

Uncompress the currently selected entries with bunzip2. The operation is performed in background (‘B-BZIP2-UNCOMPRESS’).

3.1.9.13 Encoding Files

^C e

Encode the currently selected file (‘UUENCODE’).

^C b e

Encode the currently selected file. The operation is performed in background (‘B-UUENCODE’).

^C E

Decode the currently selected file (‘UUDECODE’).

^C b E

Decode the currently selected file. The operation is performed in background (‘B-UUDECODE’).

^C k

Encode the currently selected file using mpack (‘MIME-PACK’).

^C b k

Encode the currently selected file using mpack. The operation is performed in background (‘B-MIME-PACK’).

^C K

Decode the currently selected file using munpack (‘MIME-UNPACK’).

^C b K

Decode the currently selected file using munpack. The operation is performed in background (‘B-MIME-UNPACK’).

3.1.9.14 Encrypting Files

^C p

Encrypt (using pgp) the current file (‘ENCRYPT’).

^C P

Decrypt (using pgp) the current file (‘DECRYPT’).

3.1.9.15 Comparing Files

^C =

Compare (using diff) the current ASCII file with the other panel’s current file (‘DIFF’). If both entries are directories, a recursive diff is performed.

^C ESC =

Compare (using diff) the current ASCII file with its latest backup. The latest backup is the file having the same name and a ’~’ at

the end (‘LAST-BACKUP-DIFF’).
^C B

Compare the current file with the other panel current file. A binary comparison is performed (‘compare’).

3.1.9.16 Spell Checking Files

^X I

Run the ispell command with the current file name as an argument.

3.1.9.17 Printing Files

^C j

Print the selected files via lpr to the user specified printer or to the default one if no lpr options are given (‘PRINT-JOB’).

^C J

Print the list of active printing jobs via lpq. The default printer is queried if no lpq options are given

(‘PRINT-JOB-LIST’).

3.1.9.18 Wiping Files

^C W

Call gitwipe to wipe the selected files. Asks for confirmation before actually wiping them in order to avoid errors (‘WIPE’).

See Section 3.5 [gitwipe], page 31, for more information.

3.1.9.19 Searching Files

ESC %

Search files on the file system, starting from the current directory (‘FIND’).

ESC &

Use ’locate’ to search files on the file system, starting from the current directory (‘LOCATE’).

^X w

Locate the binary, source, and manual page files for a command

(‘WHEREIS’).
^X W

Locate a command; display its pathname or alias (‘WHICH’).

3.1.9.20 Managing tar based archive files

^C a

Create a tar archive containing all the currently selected entries (‘TAR’).

^C b a

Create a tar archive containing all the currently selected entries. The operation is performed in background (‘B-TAR’).

^C x

Create a compressed tar archive containing the current entry, provided it is a directory (‘TAR-COMPRESS’).

^C b x

Create a compressed tar archive containing the current entry, provided it is directory. The operation is performed in background

(‘B-TAR-COMPRESS’).
^C -

Create a bzip2 compressed tar archive containing the current entry, provided it is a directory (‘TAR-BZIP2’).

^C b -

Create a bzip2 compressed tar archive containing the current entry, provided it is a directory. The operation is performed in background (‘B-TAR-BZIP2’).

^C X

Expand the selected archives into the current directory. The utility used is selected based on the extension of the file

(‘GENERIC-UNPACK’).
^C b X

Expand the selected archives into the current directory. The utility used is based on the extension of the file. The operation is performed in background (‘B-GENERIC-UNPACK’).

^C V

Expand the selected archives into an user supplied current directory. The utility used is selected based on the extension of the file

(‘GENERIC-UNPACK-INTO’).
^C b V

Expand the selected archives into an user supplied directory. The utility used is based on the extension of the file. The operation is performed in background (‘B-GENERIC-UNPACK-INTO’).

3.1.9.21 Working with DPKG

dpkg provides commands for manipulating and querying the dpkg database, gitfm’s DPKG commands allow quick access to most of those which operate on binary packages, files, or require package names as arguments. All commands expecting package names can also be given binary packages, the package names will be automatically extracted. Most commands use their single character dpkg option letter as the key command.

^C ^D i

Install the deb file(s) selected or pointed by the cursor

(‘DPKG-INSTALL’).
^C ^D R i

Recursively install the deb file(s) in the directories selected or pointed by the cursor (‘R-DPKG-INSTALL’).

^C ^D u

Unpack the deb file(s) selected or pointed by the cursor, but don’t configure it (‘DPKG-UNPACK’).

^C ^D R u

Recursively unpack the deb file(s) in the directories selected or pointed by the cursor, but don’t configure them (‘R-DPKG-UNPACK’).

^C ^D C

Configure the unpacked package(s) selected or pointed by the cursor

(‘DPKG-CONFIGURE’).
^C ^D r

Remove the package(s) selected or pointed by the cursor

(‘DPKG-REMOVE’).
^C ^D P

Purge the package(s) selected or pointed by the cursor

(‘DPKG-PURGE’).
^C ^D A

Update dpkg and dselect’s idea of which packages are available with information from the deb file(s) selected or pointed by the

cursor (‘DPKG-RECORDAVAIL’).
^C ^D R A

Recursively update dpkg and dselect’s idea of which packages are available with information from the deb file(s) in the directories

selected or pointed by the cursor (‘R-DPKG-RECORDAVAIL’).

^C ^D h

Display quick help file for GITFM’s DPKG commands.

(‘DPKG-HELP’).
^C ^D c

Lists the contents of the filesystem tree archive portion of the deb file pointed by the cursor (‘DPKG-CONTENTS’).

^C ^D f

Extracts control file information from a deb file pointed by the cursor (‘DPKG-FIELD’). You are presented with the file name and can either hit ENTER to see all fields, or add control file field names (space separated) to see only those fields.

^C ^D I

Provides information about a deb file pointed by the cursor

(‘DPKG-INFO’).
^C ^D l

List the package(s) selected or pointed by the cursor

(‘DPKG-LISTPKGS’).
^C ^D s

Display status details for the package(s) selected or pointed by the

cursor (‘DPKG-STATUS’).
^C ^D L

List files owned by the package(s) selected or pointed by the cursor

(‘DPKG-LISTFILES’).
^C ^D S

Find package owning file pointed to by the cursor

(‘DPKG-SEARCH’).
^C ^D p

Display available version details of package(s) selected or pointed by

the cursor (‘DPKG-PRINTAVAIL’).

3.1.9.22 Installing and Uninstalling RPM packages

Red Hat Linux Distributions use a very powerful package manager called rpm. It is used whenever you need to install/remove/upgrade/etc a software package. The packages used by rpm use the extension ‘.rpm’. gitfm provides default key bindings for some of the basic operations rpm can perform on packages: install, uninstall, upgrade and query.

^C I

Install the rpm package pointed by the cursor

(‘RPM-INSTALL’).
^C b I

Install the rpm package pointed by the cursor (‘B-RPM-INSTALL’). The operation is performed in background.

^C i

Uninstall the rpm package pointed by the cursor

(‘RPM-UNINSTALL’).
^C b i

Uninstall the rpm package pointed by the cursor (‘B-RPM-UNINSTALL’). The operation is performed in background.

^C N

Upgrade the rpm package pointed by the cursor

(‘RPM-UPGRADE’).
^C b N

Upgrade the rpm package pointed by the cursor (‘B-RPM-UPGRADE’). The operation is performed in background.

^C q

Query the rpm package manager (‘RPM-QUERY’). The default options used are ‘-qil’. See the rpm manual page for more details on using rpm.

3.1.9.23 File Types

^C t

Print the type of a file using the file utility

(‘FILE-TYPE’).

3.1.9.24 Accessing MSDOS Files

^C m d, ^C m ^D

Change the current MSDOS directory to a given directory

(‘MTOOLS-CHDIR’).
^C m P

Put the currently selected files on a MSDOS floppy

(‘MTOOLS-PUT’).
^C b m P

Put the currently selected files on a MSDOS floppy (‘B-MTOOLS-PUT’). The operation is performed in background.

^C m G

Get some files from a MSDOS floppy and put them in a given directory (‘MTOOLS-GET’).

^C b m G

Get some files from a MSDOS floppy and put them in a given directory (‘B-MTOOLS-GET’). The operation is performed in background.

^C m D

Delete files from a MSDOS floppy (‘MTOOLS-DELETE’).

^C b m D

Delete files from a MSDOS floppy (‘B-MTOOLS-DELETE’). The operation is performed in background.

^C m L

List files and directories on a MSDOS floppy (‘MTOOLS-DIR’).

^C m F
     Format a MSDOS floppy (‘MTOOLS-FORMAT’).
^C b m F

Format a MSDOS floppy (‘B-MTOOLS-FORMAT’). The operation is performed in background.

^C m M

Create a directory on a MSDOS floppy (‘MTOOLS-MKDIR’).

^C b m M

Create a directory on a MSDOS floppy (‘B-MTOOLS-MKDIR’).

The operation is performed in background.

^C m K

Remove a directory from a MSDOS floppy (‘MTOOLS-RMDIR’).

All the files and subdirectories in that directory are removed as well.

^C b m K

Remove a directory from a MSDOS floppy (‘B-MTOOLS-RMDIR’).

All the files and subdirectories in that directory are removed as well. The operation is performed in background.

^C m R

Rename a file or directory on a MSDOS floppy

     (‘MTOOLS-RENAME’).
^C b m R

Rename a file or directory on a MSDOS floppy

(‘B-MTOOLS-RENAME’). The operation is performed in background.

^C m T

Display the contents of a file located on a MSDOS floppy

     (‘MTOOLS-TYPE’).

3.1.9.25 A different action for each file type

Many files on UNIX systems have one or more extensions specifying their types. For example, a file that ends in ‘.c’ is a file containing a C program, while a file ending in ‘.tar.gz’ is a tar archive compressed with the gzip utility. Having a default action for each file type, binded on the same key, seems to be a good idea because you can use that key to obtain type specific information about a file or to process it in some type specific way much easier. The GNUIT package contains a script called gitaction that is used to detect the current file type and perform a type specific action. See Section 3.7 [gitaction], page 31, for more information.

F2, ESC 2, ^X a

Perform an action on the current file, depending on its type

(‘FILE-ACTION’).

3.1.10 Directory operations

3.1.10.1 Creating directories

F7, ESC 7, ^X M

Create a new subdirectory in the current directory with the user supplied name (‘make-directory’).

3.1.10.2 Copying directories

F5 (for directories), ESC 5 (for directories), ^C C (for directories)

Copy the currently selected entries to the user supplied path (‘copy’).

^C b C (for directories)

Copy the currently selected entries to the user supplied path. The operation is performed in background (‘B-COPY’).

3.1.10.3 Deleting directories

F8 (for subdirectories), ESC 8 (for directories), ^C D (for subdirectories)

Delete the currently selected entries (‘delete’).

^C b D (for directories)

Delete the currently selected entries. The operation is performed in background (‘B-DELETE’).

3.1.10.4 Moving directories

F6 (for directories), ESC 6 (for directories), ^C T (for directories)

Move the currently selected entries to the user supplied path (‘move’).

^C b T (for directories)

Move the currently selected entries to the user supplied path. The operation is performed in background (‘B-MOVE’).

3.1.10.5 Renaming directories

^C R (for directories)

Rename the current file or directory with the user supplied name

(‘RENAME’).
^C b R (for directories)

Rename the current file or directory with the user supplied name. The operation is performed in background (‘B-RENAME’).

3.1.10.6 Comparing Directories

^C c q

Quickly compare the files in the left panel with the files in the right one. Only the file names, sizes and time stamps are considered in the comparison.

^C c t

Compare the files in the left panel with the files in the right one. The contents of each file in the current panel will be compared against the contents of its counterpart (if any) from the other panel.

^C f d

Recursively compare (using diff -r -q) the current directory with the other panel’s current directory (‘FAST-DIFF’). Print on standard output the names of the files that differ.

^C d

Compare (using diff) the current directory with the other panel current directory. For successful operation, both panels should contain the same directory (‘DIR-DIFF’).

^C b d

Compare (using diff) the current directory with the other panel current directory. For successful operation, both panels should contain the same directory. The operation is performed in background

(‘DIR-DIFF’).

3.1.10.7 Summarize directory usage

^C U

Display the output of the du -s command on the status line

(‘DIRECTORY-USAGE’).

3.1.10.8 Changing directories

^X d, ^X ^D

Change the current working directory. The user is asked for a new directory name and the new directory is added to the directory history

(‘change-directory’).

See Section 3.1.10.9 [Dirs History], page 24, for more information.

ESC a c

Change the current directory of the current panel to the directory of

the other panel (‘adapt-current-directory’).
ESC a o

Change the current directory of the other panel to the directory of the

current panel (‘adapt-other-directory’).

3.1.10.9 Directory History

Users usually work on a limited set of subdirectories. Providing a fast method of switching between a number of intensively used directories is a good idea and gitfm has a set of builtin commands for doing it.

Usually new directories are added to the directory history when the ‘change-directory’ built-in command is used. gitfm also adds the current directory to the history list when started, when the directory history is reset and when a command having a non empty ‘new-dir’ field successfully completes its execution. See Section 4.2.3.22 [new-dir], page 38, for more information.

^X ^N

Go to the next directory in the history (‘next-directory’).

^X ^P

Go to the previous directory in the history

     (‘previous-directory’).
^X ^R

Reset the entire directory history. As explained above, the current directory becomes the only directory in the history

     (‘reset-directory-history’).

3.1.10.10 Hot Keys

gitfm provides default key bindings for switching to a number of important directories as "/", "..", "$HOME", etc.

ESC /

Go to the ‘/’ directory (‘ROOT-DIR’).

ESC .

Go to the ‘..’ directory (‘up-one-dir’).

ESC h

Go to the ‘~’ ($HOME) directory (‘HOME-DIR’).

ESC i
     Go to the ‘/usr/include’ directory (‘INCLUDE-DIR’).
ESC ESC 1
     Go to the ‘/mnt/fd0’ directory (‘FIRST-FLOPPY-DIR’).
ESC ESC 2
     Go to the ‘/mnt/fd1’ directory (‘SECOND-FLOPPY-DIR’).

3.1.11 Compiling programs

F9, ESC 9, ^X m

Run the make command in the current directory. Use -k as the default option (‘MAKE’).

^X b m

Run the make command in background in the current directory

     (‘B-MAKE’).

See Section 3.7 [gitaction], page 31, for more information.

3.1.12 Sending/receiving ascii/binary mail

^C 2 a

Send the current current ascii file by mail to an user supplied email address (‘ASCII-MAIL’).

^C b 2 a

The same as ‘ASCII-MAIL’, the only difference being that the command runs in background (‘B-ASCII-MAIL’).

^C 2 b

Send the current current binary file by mail to an user supplied list of email addresses. The file is uuencoded first (‘BINARY-MAIL’).

^C b 2 b

The same as ‘BINARY-MAIL’, the only difference being that the command runs in background (‘B-BINARY-MAIL’).

^C 2 m

Send the current current binary file by mail to an user supplied list of email addresses. The file is encoded with mpack first

(‘MIME-MAIL’).
^C b 2 m

The same as ‘MIME-MAIL’, the only difference being that the command runs in background (‘B-MIME-MAIL’).

ESC x r m

Run the emacs -f rmail command. This will start the Emacs’s ‘rmail’ function so that you can read your mail (‘READ-MAIL’).

3.1.13 Starting a sub-shell

^X z

Call a sub-shell as specified by the $GNUIT SHELL environment variable (‘SUB-SHELL’).

$GNUIT SHELL used to be called $GIT SHELL. The old name is still accepted for backwards-compatibility.

See Section 4.1 [Environment Variables], page 33, for more information.

3.1.14 Using grep and recursive grep

^X g

Search using grep all the selected files for a given pattern (‘GREP’).

^X g

Search recursively using gitrgrep all the user specified files and directories for a given pattern (‘RECURSIVE-GREP’).

See Section 3.9 [gitrgrep], page 32, for more information.

3.1.15 Locking your console

Having a lock feature might be a good idea and, since not all the UNIX systems provide one, gitfm tries to get around the problem …

^X p

Prompt the user for a password and locks the console until the same password is reinserted (‘lock’).

3.1.16 Refreshing the screen contents

Sometimes your screen needs to be refreshed. Just think about what happens when somebody wants to talk with you and the talk daemon writes something like this

Message from Talk Daemon@galei.cs.vu.nl at 12:15 ... talk: connection requested by andrei@galei.cs.vu.nl. talk: respond with: talk andrei@galei.cs.vu.nl

on your screen. And sometimes you might also want to re-read the current directories. gitfm provides a built-in command for refreshing the screen contents.

^L

Re-read the directories contents and refresh the screen (‘refresh’).

3.1.17 Resetting your terminal

^X ^L

Call reset in order to reset the terminal to its default

settings (‘TTY-RESET’).

3.1.18 Mounting/unmounting file systems

People dealing with lots of files usually need to save/restore/copy files from/to other file systems. In order to be more efficient, gitfm provides a set of key bindings for mounting and unmounting file systems. See Section 3.6 [gitmount], page 31, for more information.

The default key bindings set has been designed to work under Linux, but it can be easily changed for other UNIX systems with different device names. Reading the configuration file ‘gnuitrc.common’ should be enough. See Section 3.1.10.10 [Hot Keys], page 24, for more information.

As a convention, the ‘/mnt’ directory is used to store an empty subdirectory for each mountable file system. Each file system is actually mounted in its counterpart ‘/mnt’ subdirectory. Try to follow this convention since the gitmount script is heavily based on it. See Chapter 4 [Customization], page 33, for more information.

ESC m a

Call mount(1) in order to mount the first floppy

(‘/dev/fd0’) in the ‘/mnt/fd0’ directory (‘MOUNT-A’).

ESC m b

Call mount(1) in order to mount the second floppy (‘/dev/fd1’) in the ‘/mnt/fd1’ directory (‘MOUNT-B’).

ESC m c

Call mount(1) in order to mount the cdrom

(‘/dev/cdrom’) in the ‘/mnt/cdrom’ directory (‘MOUNT-CDROM’).
ESC m f

Call mount(1) in order to mount the first floppy

(‘/dev/fd0’) in the ‘/mnt/floppy’ directory (‘MOUNT-FLOPPY’).
ESC m z

Call mount(1) in order to mount the zip drive

(‘/dev/zip’) in the ‘/mnt/zip’ directory (‘MOUNT-ZIP’).
ESC m j

Call mount(1) in order to mount the jaz drive

(‘/dev/jaz’) in the ‘/mnt/jaz’ directory (‘MOUNT-JAZ’).
ESC m t

Call mount(1) in order to mount the file systems corresponding to the selected subdirectories. For example, if you are in the ‘/mnt’ directory and the ‘cdrom’ and ‘zip’ subdirectories are selected, the cdrom and the zip disk will be mounted (‘MOUNT-THESE’).

ESC r a

Call umount(1) in order to remove (unmount) the first floppy

(‘/dev/fd0’) (‘UMOUNT-A’).
ESC r b

Call umount(1) in order to remove (unmount) the second floppy

(‘/dev/fd1’) (‘UMOUNT-B’).
ESC r c

Call umount(1) in order to remove (unmount) the cdrom

(‘/dev/cdrom’) (‘UMOUNT-CDROM’).

findex UMOUNT-CDROM

ESC r f

Call umount(1) in order to remove (unmount) the first floppy

(‘/dev/fd0’) (‘UMOUNT-FLOPPY’).
ESC r z

Call umount(1) in order to remove (unmount) the zip drive

(‘/dev/zip’) (‘UMOUNT-ZIP’).
ESC r j

Call umount(1) in order to remove (unmount) the jaz drive

(‘/dev/jaz’) (‘UMOUNT-JAZ’).
ESC r t

Call umount(1) in order to remove (unmount) the file systems mounted into the selected subdirectories. For example, if the current directory is ‘/mnt’ and the ‘cdrom’ and ‘zip’ subdirectories are selected, the cdrom and the zip disk will be unmounted (‘UMOUNT-THESE’).

3.1.19 Getting some useful system information

^X T

Call date(1) in order to display the current time/date (‘DATE’).

ESC S f

Call finger(1) in order to display information about local and remote users (‘FINGER’).

ESC S m

Call mount(1) in order to display a list of the currently mounted file systems (‘MOUNTED-FILE-SYSTEMS’).

ESC S q

Call quota(1) in order to display a user file system disk quota and quota (‘QUOTA’).

ESC S s

Call df(1) in order to get the status of the currently mounted

file systems (‘DISK-FREE-SPACE’).
ESC S u

Call users(1) in order to get the name of the currently logged in users (‘USERS’).

ESC S v

Call $GNUIT_VMSTAT(1) in order to get the current virtual memory status. This is very system dependent, Linux uses free, other systems use vmstat, so the $GNUIT VMSTAT variable is used to deal with this (‘VIRTUAL-MEMORY-STATUS’).

$GNUIT VMSTAT used to be called $GIT VMSTAT. The old name is still accepted for backwards-compatibility.

See Section 4.1 [Environment Variables], page 33, for more information.

ESC S w

Call who(1) in order to find out who is on the system (‘WHO’).

3.1.20 How to look at the environment variables

^X E

Call env(1) in order to display the current environment (‘ENV’).

^X H

Call xhost(1) in order to add/remove hosts names to the list allowed to make connection to the X server (‘XHOST’).

3.1.21 Viewing/killing processes

There are at least two kinds of ps(1) utilities. One that accepts (more or less) combinations of the ’a’, ’u’, and ’x’ flags and another that accepts combinations of ’e’, ’f’ and ’l’ flags. Since is quite difficult to test which one works fine on a given UNIX system, gitfm provides key bindings for both of them. Anyway, if your ps(1) fails to accept the predefined combinations, please take a look in its manual and then modify the ‘.gnuitrc.TERM’ file as needed.

Since the number of possible combinations of flags in the ps command line is quite big and *very* system dependent, there is no real reason to display them all here. We are only interested in giving you a starting point in your search through the ‘.gnuitrc.TERM’ file.

Note also that you can display a list of processes using ps(1) or browse through a list of them (killing as needed) using gitps. As a convention, we have used the same key sequence for a given set of ps(1) flags for both ps(1) and gitps, the only difference being that ps(1) keys end in an uppercase letter. See Section 3.2 [gitps], page 30, for more information.

Under Linux it is possible to see a tree of processes using pstree(1).

Here there are the default key bindings for the ’e’, ’f’ and ’l’ ps(1) flags combinations:

ESC P b, ESC P c, ESC P e

Call gitps or ps(1) in order to browse through or display

a list of currently running processes (‘GITPS’, ‘PS’). ... and the default key bindings for the ’a’, ’u’ and ’x’ ps(1) flags combinations:

ESC P a, ESC P l, ESC P u ESC P x, ESC P y

Call gitps or ps(1) in order to browse through or display

a list of currently running processes (‘GITPS’, ‘PS’).

ESC P T

Call pstree(1) in order to display the tree of currently running

processes (‘PSTREE’).

^X k

Call kill(1) in order to kill a user specified process with a

given signal (‘KILL’).

3.1.22 Synchronizing the file systems

^X S

Call sync(1) in order to synchronize all the file systems

(‘SYNC’).

3.1.23 Reading the documentation

^X q

Read a manual page. The user is prompted for its name (‘MAN’).

F1, ESC 1, ^X i

Read an info documentation. The user is prompted for the documentation name (‘INFO’).

^X h

Read the html documentation using the viewer specified in GNUIT BROWSER, or with lynx if GNUIT BROWSER is not set (‘HTML’).

$GNUIT BROWSER used to be called $GIT BROWSER. The old name is still accepted for backwards-compatibility.

^X ^A

Show the key-bindings for commands matching a specified string

     (‘apropos’).

3.1.24 Exiting GNU Interactive Tools

F10, ESC 0, ^X ^C, ^X c

Exit GNU Interactive Tools (‘exit’).

3.2 The GNUIT process viewer/killer

gitps is an interactive process viewer/killer. It calls internally the ps(1) utility. This is a brief description of the command line arguments. -h print this help message -v print the version number -i print the installation directory -c use ANSI colors -b don’t use ANSI colors -l don’t use the last screen character -p pass the remaining arguments to ps(1)

Running gitps is self explanatory. Use the arrows, PageUp, PageDown, Home, End, ^N, ^P, ^V, ESC v, Space and Backspace to move in the list, ^L to refresh it, Enter to change the default signal and F10, q or ^X ^C to leave.

You can change these keys, just read the GITPS-Setup, GITPS-Color, GITPS-Monochrome and GITPS-Keys sections in the configuration files ‘gnuitrc.TERM’.

The selected signal can also be changed by pressing its first letter (in uppercase), e.g. for ‘SIGTERM’ press T, or by pressing the key corresponding to the signal number. For signals > 10, prefix with ^X, and for signals greater than 20, prefix with ^C. For instance, to select signal 15 (‘SIGTERM’), press ^X5.

Keys for signal numbers are hard-coded to the versions in i386 Linux, however most correspond to POSIX. Compare signal(7) or signal(5) with

http://linux.die.net/man/7/signal.

3.3 The GNUIT ASCII/HEX file viewer

gitview is an ASCII/HEX file viewer. Use the arrows, PageUp, PageDown, Home, End, ^N,

^P, ^V, ESC v, Space and Backspace to move in the file, ^L to refresh the screen and F10, q or ^X ^C to leave.

You can change these keys, just read the GITVIEW-Setup, GITVIEW-Color, GITVIEW-Monochrome and GITVIEW-Keys sections in the configuration files

‘gnuitrc.TERM’.

Here is a brief description of the command line arguments: -h print this help message -v print the version number -i print the installation directory -c use ANSI colors -b don’t use ANSI colors -l don’t use the last screen character

3.4 The GNUIT key sequences display utility

gitkeys is a program that displays the key sequence sent by the pressed key. This is the key sequence received by GIT tools, so this program is useful when setting up the ‘.gnuitrc.TERM’ configuration files.

3.5 The GNUIT wipe file utility

gitwipe is an utility for wiping files. It overwrites the file contents with a random sequence of numbers and then calls ‘sync’().

Note that gitwipe does *not* remove the wiped file since (under Linux at least) the ‘sync’() system call might return before actually writing the new file contents to disk. Removing the file might be dangerous because some file systems can detect that the blocks in the removed wiped file are no longer used and never write them back to disk in order to improve performance. It is up to you to remove the file(s) at a later moment.

3.6 The GNUIT mount utility

gitmount is a script that allows you to mount a list of block devices (specified in the command line), without specifying the file system type. With a command like ‘gitmount fd0 cdrom’ the first floppy will be mounted in ‘/mnt/fd0’ and the cdrom will be mounted in ‘/mnt/cdrom’. Make sure your ‘/etc/fstab’ settings are correct.

You don’t need to know the file system type anymore. If you want to use gitmount with the block device ‘/dev/xxx’ then the directory ‘/mnt/xxx’ is created if it doesn’t exist. gitmount will attempt to create the necessary directories, but root permissions might be required.

3.7 The GNUIT per file type action script

gitaction is a script that executes a different action for each file type specified. It is called by the gitfm program when pressing F2, ESC 2 or ^Xa.

The first parameter is the current directory name and the second one is the file name to be matched against the default patterns. The matching is done using the shell ’case’ statement.

If you press F2, ESC 2 or ^Xa on a ‘*.html’ file, gitfm will invoke a browser to view it, if you press F2, ESC 2 or ^Xa on a ‘*.tar.gz’ file, gitfm will list the tar archive contents, if you press the same keys on a ‘*.gz’ file, gitfm will display its uncompressed contents on the screen, etc …

If you press F2, ESC 2 or ^Xa on a ‘*.gif’ file or ‘*.jpg’ file and you have the xzgv utility installed, you will be able to see it. If you want to change the gif/jpeg viewer, all you need to do is to change its name in the gitaction script.

There are many more file types and viewers that gitaction knows about. In addition, if you are running under GNOME, gnome-open is used, and on MacOS, open(1) is used. see(1) and metamail(1) are also used as fallbacks. If all else fails, the file is displayed using $GNUIT PAGER.

If you want to find out what the default action for each file type is (or if you want to modify it), just read/modify the gitaction script.

Also, you can add a .gitaction shell script in your home directory and/or in any other directory. Before trying to match a file name, gitaction will attempt to execute ./.gitaction. If that one fails to match the file name against its patterns, it backs up to $HOME/.gitaction. When this one fails too the patterns in gitaction are tried. For an example of how to write .gitaction scripts take a look at the .gitaction shell script provided as part of the distribution and installed in the ‘$(prefix)/bin’ directory. Debian

users will find it in ‘/usr/share/doc/gnuit/examples/’.

3.8 Unified archive unpacking

gitunpack is a shell script that accepts a directory and a set of archives as its command line parameters, and then attempts to unpack those archives in the given directory, selecting the utility used to unpack the archives based on the archive extensions.

3.9 The GIT recursive grep script

gitrgrep is a very small script that calls grep recursively. It accepts grep like options / parameters, the only difference being that file specifications should be quoted:

     gitrgrep main '*.c'
     or
     gitrgrep errno '*.c *.h'

gitregrep and gitrfgrep are recursive versions of the egrep and fgrep programs.

4 Customizing GNU Interactive Tools

4.1 Environment Variables

The configuration files use shell environment variables to call the shell, editor, mail reader, html viewer, compress and virtual memory status utility. That means that if you set GNUIT SHELL, GNUIT EDITOR, GNUIT RMAIL, GNUIT BROWSER, or GNUIT VMSTAT to some value, that value will be used instead of the default one. The defaults are:

GNUIT SHELL=’/bin/sh’ GNUIT EDITOR=’vi’ GNUIT RMAIL=’emacs -f rmail’ GNUIT PAGER=’more GNUIT VMSTAT=’free’ GNUIT BROWSER=’lynx’

if the configure script is passed ‘--enable-debian’, some of the defaults are changed as follows:

GNUIT EDITOR=’sensible-editor’ GNUIT PAGER=’sensible-pager’ GNUIT BROWSER=’sensible-browser’

If SHELL is defined, GNUIT SHELL will be set to that value. If PAGER is defined, GNUIT PAGER will be set to that value. If EDITOR is defined, GNUIT EDITOR will be set to that value. If you want to change the default settings, put something like this into

your ‘.profile’:

export GNUIT SHELL=’/usr/local/bin/bash’ export GNUIT EDITOR=’emacs’ export GNUIT RMAIL=’elm’ export GNUIT PAGER=’less’ export GNUIT VMSTAT=’vmstat’ export GNUIT BROWSER=’netscape’

These variables used to be prefixed with GIT (e.g. GIT PAGER). The old names are still accepted for backwards compatibility.

4.2 Configuration Files

There is one configuration file per terminal type in GNUIT. The configuration file(s) reside in the user’s home directory or (the default versions) in the directory ‘/usr/share/gnuit/’). Files in the user’s home directory start with a leading “.”

Their generic name is ‘gnuitrc.TERM’. GNUIT allows each terminal type to have its own configuration file (TERM is the value of the TERM environment variable (e.g ‘vt102’); for the Linux console the configuration file is ‘gnuitrc.console’).

Since most of the key bindings are common to all the terminal types, a configuration file called ‘gnuitrc.common’ is parsed before parsing the normal ‘.gnuitrc.TERM’ configuration file, the later one defining only those keys that are terminal specific. However, if a key binding is redefined in the ‘gnuitrc.TERM’ file, that binding will be used.

If the GNUIT package have been compiled without passing the ‘--enable-terminfo’ option to the configure script and your system has a huge ‘termcap’ database (‘/etc/termcap’), you can copy the termcap definition(s) of your terminal(s) in a file called, lets say ‘.termcap’ and put it in your home directory. After that, set your TERMCAP environment variable to point to it. You should add something like this to

your ‘.profile’:
     TERMCAP=‘/home/mike/.termcap’

The interactive programs in the GNUIT package can run without such a file, but on systems with huge ‘termcap’ databases, copying the definitions of the most used terminals in a local ‘.termcap’ file will lead to a faster start.

The ‘.gnuitrc.TERM’ is first searched for in the home directory then, if not found, in the directory ‘/usr/share/gnuit/’). (without the leading “.”). The configuration file is structured in sections, each section containing variables in the following format:

     ‘variable-name’ = ‘first-field’;‘second-field’; ...

After the ‘variable-name’ at least one space or tab is required. All characters after a ‘#’ are ignored and if you comment a section name, the whole section is ignored.

Section names are enclosed in rectangular brackets (‘[’ and ‘]’). Note that this manual don’t include them while refering to section names.

The GNUIT package contains three major programs: gitfm, gitps and gitview. Each one has its own sections in the configuration files. There is also a global setup section called ‘Setup’ that is used by all these programs.

Note that now git has been renamed to gitfm, the corresponding sections have been renamed to GITFM, however for backwards compatibility, the old name (GIT) is still checked if the GITFM section is not found.

4.2.1 Writing key sequences

GNUIT contains three interactive programs. Their names are: gitfm (this is the file system browser), gitps (this is the process viewer/killer and gitview (this is the ASCII/HEX file viewer). Each one of these programs has its own set of key bindings.

The convention used in describing key bindings are very simple. Here there are some examples that will help you to understand them. The corresponding Emacs conventions will help you even more.

^A means keeping the Ctrl key down and pressing the a key (C-a).

The ESC character is represented as ^[ so that you can use the meta character (M- ) where available (or the ESC key):

^[a corresponds to M-a (pressing the ESC key and then a).

The ^character is represented as ^^.

The backspace character is represented as ^_.

The Ctrl-SPACE character (C-SPC) is represented as ^$.

The space (SPC) character is represented as ^@.

Note that the key bindings notation described here is only used in the configuration files. For the sake of readability this manual uses ESC for the ESC key, SPC for the SPACE key and

RET for the RETURN (ENTER) key.

4.2.2 The global setup section

In this section the variables have only one field.

‘AnsiColors’

This variable should be set to ‘ON’ if the terminal supports standard ‘ANSI’ color sequences. Otherwise it should be ‘OFF’. If ‘AnsiColors’ is ‘ON’, ‘GITxxx-Color’ sections will be used in the configuration files ‘gnuitrc.TERM’. Otherwise, GNUIT interactive programs will use the ‘GITxxx-Monochrome’ sections.

‘UseLastScreenChar’

This variable is used for terminals that can’t write on the last character of the screen without scrolling the entire screen. If your terminal has no problem writing there (Linux console, vt100, vt102, xterm, ...) set it to ‘ON’. Otherwise (hpterm), it should be ‘OFF’.

‘StartupScrollStep’

This variable specifies the scroll step initial value for both panels.

4.2.3 gitfm Sections

4.2.3.1 gitfm Setup

In this section the variables have only one field.

‘StartupFileDisplayMode’

This variable specifies the file specific information displayed at startup. It can be any

of ‘OwnerGroup’, ‘DateTime’, ‘Size’, ‘AbbrevSize’, ‘Mode’ or ‘FullName’. Its value initially

affects both panels but it can be changed separately afterward.

‘StartupFileSortMethod’

This variable specifies the startup sort method. It can be any of ‘Name’, ‘Extension’,

‘Size’, ‘Date’, ‘Mode’, ‘OwnerId’, ‘GroupId’, ‘OwnerName’ or ‘GroupName’. Its value initially

affects both panels but it can be changed separately afterward.

‘MaxUnscaledDigits’

Maximum number of digits a number may be before being scaled (e.g. to ‘123M’). If you want number to always be scaled, either set this to ‘0’, or set ‘StartupFileDisplayMode’ to ‘AbbrevSize’. Note that scaling may happen anyway if the number is too large for the display field.

‘GroupDigits’

If this variable is ‘ON’, digits of file sizes will be grouped according to your locale, (e.g.

‘123,456,789’).
‘ConfirmOnExit’

If this variable is ‘ON’, the user is prompted for confirmation at exit.

‘HistoryFile’

This variable specifies the history file name. The default value is ‘~/.githistory’.

‘InfoDisplay’

If this variable is ‘OFF’, auxiliary file informations are not displayed. This can be useful if you are using a very slow terminal.

‘LeadingDotMatch’

If this variable is ‘OFF’ when matching files for select-files-matching-pattern / unselectfiles-matching-pattern then the leading ’.’ in the file name is matched only explicitly.

‘TypeSensitivity’

If this variable is ‘OFF’, colors are not used when displaying files. Normally, the information in the ‘GITFM-FTI’ section is used to display files with different colors, depending on their types. Note that ‘TypeSensitivity’ is automatically set to ‘OFF’ when ‘AnsiColors’ is ‘OFF’. See Section 4.2.4 [GITFM-FTI], page 39, for mor information.

‘NormalModeHelp’ ‘CommandLineModeHelp’

These variables describe the status bar contents for each gitfm mode when no errors occurred. gitfm can display on the status bar a help string and/or some system information (system type, hostname, machine type and the current date) using escape characters:

\s -> the system type \h -> the host name \m -> the machine type \d -> the current date

See Section 3.1.3 [Modes], page 4, for more information.

4.2.3.2 Using gitfm on color displays

In this sections the variables have only one field.

These section allows you to customize the colors of gitfm. Reading the ‘gnuitrc.TERM’ configuration file is self explanatory.

4.2.3.3 Using gitfm on monochrome displays

In this sections the variables have only one field.

These section allows you to customize the appearance of gitfm on monochrome displays. Reading the ‘gnuitrc.TERM’ configuration file is self explanatory.

4.2.3.4 Defining keys

These section describes the actions gitfm takes when a specified key is pressed. A variable can have up to 6 fields separated by ’;’. Each line in this section looks like:

     ‘key-sequence’ = ‘command-name’;‘formatted-command’;‘new-dir’;
         ‘save-screen’;‘pause’;‘hide’

Note that you can’t continue the variable fields description on the next line.

4.2.3.5 The key-sequence field

‘key-sequence’ is the key sequence associated with the given command. You can use any key sequence that doesn’t start with an ascii character (0x20 to 0x7e).

   Symbolic key names (F0, F1, F2, ... F10, UP, DOWN, RIGHT, LEFT, INS, DEL, HOME, END,

PGUP and PGDOWN) can be used instead of the key sequence. If some keys don’t have a ‘termcap’/ ‘terminfo’ description (like the F11/F12 keys on the Linux console) you can specify the key sequence in the usual way.

4.2.3.6 The command-name field

‘command-name’ is a command generic name. Even if it is not always used, the ‘command-name’ must be present (if a command is associated with a ‘key-sequence’). If it is not, no action will be taken when pressing ‘key-sequence’.

There are two types of commands in gitfm: built-in commands and user defined commands. If the ‘command-name’ section contains a built-in command specification, the other fields are ignored.

Note that by convention built-in command names contain only lower case letters while user defined command names contain only upper case letters.

4.2.3.7 The formatted-command field

‘formatted-command’ is a shell command which can contain some scanf(3)-like format specifiers. They are used to get the current entry name, owner, group, mode, etc.

Note that using uppercase ‘format specifiers’ you will be able to access the other panel path, file and directory names, etc.

These are the available ‘format specifiers’:

4.2.3.8 The %s format specifier

The format of %s is: %s{question,default answer}.

When gitfm encounters a %s in the ‘formatted-command’ it asks the user the question ‘question’ whose default answer is ‘default_answer’ and replaces the ‘%s{ , }’ with the user’s answer. Both ‘question’ and ‘default_answer’ can contain any other ‘format

specifiers’ except %s.

Note that there should be no spaces between %s and ’{’.

4.2.3.9 The %f format specifier

gitfm will replace %f with the current directory entry name only if it is a file (not a directory).

4.2.3.10 The %d format specifier

gitfm will replace %d with the current directory entry name only if it is a directory (not a file).

4.2.3.11 The %l format specifier

gitfm will replace %l with the current directory entry name only if it is a symbolic link with no target.

4.2.3.12 The %t format specifier

gitfm will replace %t with the current directory entry name only if it is a named pipe.

4.2.3.13 The %z format specifier

gitfm will replace %z with the current directory entry name only if it is a socket.

4.2.3.14 The %a format specifier

gitfm will always replace %a with the current directory entry name.

4.2.3.15 The %m format specifier

gitfm will always replace %m with the current file mode.

4.2.3.16 The %g format specifier

gitfm will always replace %g with the current file group.

4.2.3.17 The %o format specifier

gitfm will always replace %o with the current file owner.

4.2.3.18 The %p format specifier

gitfm will always replace %p with the current panel path.

4.2.3.19 The %b format specifier

gitfm will always replace %b with the current panel directory name.

4.2.3.20 The %i format specifier

gitfm will always replace %i with all the current panel selected entry names.

4.2.3.21 The %? format specifier

The format of %? is: %?{confirmation}.

gitfm uses this format specifier only to ask for confirmation before expanding / executing the current command. The ‘confirmation’ string is displayed and, if the user doesn’t confirm, the command is aborted. Otherwise, %?{confirmation} expands to a null string and the command is expanded / executed normally.

4.2.3.22 The new-dir field

If the ‘formatted-command’ successfully exits (exit code = 0) or it has no body and this field is present then ‘new-dir’ will become the current panel directory.

The character ’~’ used at the beginning of the ‘new-dir’ field is replaced by the user’s home directory.

4.2.3.23 The save-screen field

This field is a character (usually ’y’ or ’n’) that tells gitfm to save (’y’) or not to save (’n’) the terminal’s screen after executing the ‘formatted-command’. Saving the screen is not necessary while editing or viewing a file because the information left after the editor or the viewer exits is not important. Saving the screen means that that screen will be restored before the execution of the next command. Currently this field is used only if you are working as a super user under Linux on a virtual console. Its default value is ’y’.

4.2.3.24 The pause field

Users may wish to read the result of some commands before repainting the panels. If this field is present gitfm will wait for a key to be pressed before restoring the panels. Its default value is ’n’.

4.2.3.25 The hide field

Some commands that don’t displaying any useful information if successfully complete their execution: mount, chmod, chown, chgrp, sync ... and, if an error occurs, a line or two are sent to stderr. If this option is ’y’, the stdout and stderr will be redirected to some files (‘git.1.pid’ and ‘git.2.pid’, where pid is gitfm’s pid) and only if the command’s exit code is not 0, the ‘git.2.pid’ file will be displayed, line by line, onto the status bar. This way the panels will not be deleted and then repainted and the command appears to be built-in. ‘git.1.pid’ and ‘git.2.pid’ are created in the temporary directory specified in the TMPDIR environment variable (or "/tmp" if TMPDIR is not defined). The default value of the hide field is ’n’.

4.2.4 Setting up colors for different file types

This sections contains entries of the form:

     `pattern' = `foreground'; `background'; `brightness'

where ‘pattern’ is a file name matching pattern, ‘foreground’, ‘background’ and ‘brightness’ are the color specification to be used when a file whose name match the given ‘pattern’ is displayed in a panel. Colors can be turned off using the ‘TypeSensitivity’ variable in the ‘GITFM-Setup’ section.

4.2.5 gitps Sections

4.2.5.1 gitps Setup

In this section the variables have only one field.

‘Help’

This variable describe gitps’s status bar contents.

4.2.5.2 Using gitps on color displays

In this sections the variables have only one field.

These section allows you to customize the colors of gitps. Reading the ‘gnuitrc.TERM’ configuration file is self explanatory.

4.2.5.3 Using gitps on monochrome displays

In this sections the variables have only one field.

These section allows you to customize the appearance of gitps on monochrome displays. Reading the ‘gnuitrc.TERM’ configuration file is self explanatory.

4.2.5.4 Defining keys

4.2.6 gitview Sections

4.2.6.1 gitview Setup

In this section the variables have only one field.

‘Help’

This variable describe gitps’s status bar contents.

4.2.6.2 Using gitview on color displays

In this sections the variables have only one field.

These section allows you to customize the colors of gitview. Reading the ‘gnuitrc.TERM’ configuration file is self explanatory.

4.2.6.3 Using gitview on monochrome displays

In this sections the variables have only one field.

These section allows you to customize the appearance of gitview on monochrome displays. Reading the ‘gnuitrc.TERM’ configuration file is self explanatory.

4.2.6.4 Defining keys

5 GNU Interactive Tools limitations

Background commands (& terminated)can be specified in the configuration file but their result (stdout and stderr redirection), will be overwritten by the result of newer commands and, if an error occurs, it will not be seen.

When gitfm is compiled for Linux, the default built-in color descriptions are for color monitors, so you can’t (decently) run gitfm on a b/w monitor without the ‘gnuitrc.TERM’ file correctly configured. ‘gnuitrc.TERM’ should be configured with ‘AnsiColors’ = OFF. However, if your system knows about the linux-m terminal type, using that may be a better solution.

Job support is implemented only in gitfm.

Due to the fact that the ’;’ character is used as a field separator in the configuration files, you can’t write something like that in the ‘gnuitrc.TERM’ files:

^AAA = SHOW-USERS-AND-GROUPS; more /etc/passwd; more /etc/group because ’more /etc/group’ will be considered as a directory to switch to. You must write a small script instead:

#! /bin/sh

more /etc/passwd more /etc/group

Supposing the script name is show_ug, the ‘gnuitrc.TERM’ line will look like this:

^AAA = SHOW-USERS-AND-GROUPS; show ug

There is no support for appearance modes on magic-cookie terminals.

Chapter 6: GNU Interactive Tools bugs 42

6 GNU Interactive Tools bugs

Any questions, comments, or bug reports, should be emailed to ‘gnuit-dev@gnu.org’. Please include the version number.

Appendix A Copying This Manual

Version 1.3, 3 November 2008

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2000, 2001, 2002, 2007, 2008 Free Software Foundation, Inc.

 http://fsf.org/

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

0. PREAMBLE The purpose of this License is to make a manual, textbook, or other functional and useful document free in the sense of freedom: to assure everyone the effective freedom to copy and redistribute it, with or without modifying it, either commercially or noncommercially. Secondarily, this License preserves for the author and publisher a way to get credit for their work, while not being considered responsible for modifications made by others. This License is a kind of “copyleft”, which means that derivative works of the document must themselves be free in the same sense. It complements the GNU General Public License, which is a copyleft license designed for free software. We have designed this License in order to use it for manuals for free software, because free software needs free documentation: a free program should come with manuals providing the same freedoms that the software does. But this License is not limited to software manuals; it can be used for any textual work, regardless of subject matter or whether it is published as a printed book. We recommend this License principally for works whose purpose is instruction or reference.

1. APPLICABILITY AND DEFINITIONS This License applies to any manual or other work, in any medium, that contains a notice placed by the copyright holder saying it can be distributed under the terms of this License. Such a notice grants a world-wide, royalty-free license, unlimited in duration, to use that work under the conditions stated herein. The “Document”, below, refers to any such manual or work. Any member of the public is a licensee, and is addressed as “you”. You accept the license if you copy, modify or distribute the work in a way requiring permission under copyright law. A “Modified Version” of the Document means any work containing the Document or a portion of it, either copied verbatim, or with modifications and/or translated into another language. A “Secondary Section” is a named appendix or a front-matter section of the Document that deals exclusively with the relationship of the publishers or authors of the Document to the Document’s overall subject (or to related matters) and contains nothing that could fall directly within that overall subject. (Thus, if the Document is in part a textbook of mathematics, a Secondary Section may not explain any mathematics.) The relationship could be a matter of historical connection with the subject or with related matters, or of legal, commercial, philosophical, ethical or political position regarding them. The “Invariant Sections” are certain Secondary Sections whose titles are designated, as being those of Invariant Sections, in the notice that says that the Document is released under this License. If a section does not fit the above definition of Secondary then it is not allowed to be designated as Invariant. The Document may contain zero Invariant Sections. If the Document does not identify any Invariant Sections then there are none.

The “Cover Texts” are certain short passages of text that are listed, as Front-Cover Texts or Back-Cover Texts, in the notice that says that the Document is released under this License. A Front-Cover Text may be at most 5 words, and a Back-Cover Text may be at most 25 words.

A “Transparent” copy of the Document means a machine-readable copy, represented in a format whose specification is available to the general public, that is suitable for revising the document straightforwardly with generic text editors or (for images composed of pixels) generic paint programs or (for drawings) some widely available drawing editor, and that is suitable for input to text formatters or for automatic translation to a variety of formats suitable for input to text formatters. A copy made in an otherwise Transparent file format whose markup, or absence of markup, has been arranged to thwart or discourage subsequent modification by readers is not Transparent. An image format is not Transparent if used for any substantial amount of text. A copy that is not “Transparent” is called “Opaque”.

Examples of suitable formats for Transparent copies include plain ascii without markup, Texinfo input format, LaTEX input format, SGML or XML using a publicly available DTD, and standard-conforming simple HTML, PostScript or PDF designed for human modification. Examples of transparent image formats include PNG, XCF and JPG. Opaque formats include proprietary formats that can be read and edited only by proprietary word processors, SGML or XML for which the DTD and/or processing tools are not generally available, and the machine-generated HTML, PostScript or PDF produced by some word processors for output purposes only.

The “Title Page” means, for a printed book, the title page itself, plus such following pages as are needed to hold, legibly, the material this License requires to appear in the title page. For works in formats which do not have any title page as such, “Title Page” means the text near the most prominent appearance of the work’s title, preceding the beginning of the body of the text.

The “publisher” means any person or entity that distributes copies of the Document to the public.

A section “Entitled XYZ” means a named subunit of the Document whose title either is precisely XYZ or contains XYZ in parentheses following text that translates XYZ in another language. (Here XYZ stands for a specific section name mentioned below, such as “Acknowledgements”, “Dedications”, “Endorsements”, or “History”.) To “Preserve the Title” of such a section when you modify the Document means that it remains a section “Entitled XYZ” according to this definition.

The Document may include Warranty Disclaimers next to the notice which states that this License applies to the Document. These Warranty Disclaimers are considered to be included by reference in this License, but only as regards disclaiming warranties: any other implication that these Warranty Disclaimers may have is void and has no effect on the meaning of this License.

2. VERBATIM COPYING You may copy and distribute the Document in any medium, either commercially or noncommercially, provided that this License, the copyright notices, and the license notice saying this License applies to the Document are reproduced in all copies, and that you add no other conditions whatsoever to those of this License. You may not use technical measures to obstruct or control the reading or further copying of the copies you make or distribute. However, you may accept compensation in exchange for copies. If you distribute a large enough number of copies you must also follow the conditions in section 3. You may also lend copies, under the same conditions stated above, and you may publicly display copies.

3. COPYING IN QUANTITY If you publish printed copies (or copies in media that commonly have printed covers) of the Document, numbering more than 100, and the Document’s license notice requires Cover Texts, you must enclose the copies in covers that carry, clearly and legibly, all these Cover Texts: Front-Cover Texts on the front cover, and Back-Cover Texts on the back cover. Both covers must also clearly and legibly identify you as the publisher of these copies. The front cover must present the full title with all words of the title equally prominent and visible. You may add other material on the covers in addition. Copying with changes limited to the covers, as long as they preserve the title of the Document and satisfy these conditions, can be treated as verbatim copying in other respects. If the required texts for either cover are too voluminous to fit legibly, you should put the first ones listed (as many as fit reasonably) on the actual cover, and continue the rest onto adjacent pages. If you publish or distribute Opaque copies of the Document numbering more than 100, you must either include a machine-readable Transparent copy along with each Opaque copy, or state in or with each Opaque copy a computer-network location from which the general network-using public has access to download using public-standard network protocols a complete Transparent copy of the Document, free of added material. If you use the latter option, you must take reasonably prudent steps, when you begin distribution of Opaque copies in quantity, to ensure that this Transparent copy will remain thus accessible at the stated location until at least one year after the last time you distribute an Opaque copy (directly or through your agents or retailers) of that edition to the public. It is requested, but not required, that you contact the authors of the Document well before redistributing any large number of copies, to give them a chance to provide you with an updated version of the Document.

4. MODIFICATIONS You may copy and distribute a Modified Version of the Document under the conditions of sections 2 and 3 above, provided that you release the Modified Version under precisely this License, with the Modified Version filling the role of the Document, thus licensing distribution and modification of the Modified Version to whoever possesses a copy of it. In addition, you must do these things in the Modified Version: A. Use in the Title Page (and on the covers, if any) a title distinct from that of the

Document, and from those of previous versions (which should, if there were any, be listed in the History section of the Document). You may use the same title as a previous version if the original publisher of that version gives permission.

B. List on the Title Page, as authors, one or more persons or entities responsible for authorship of the modifications in the Modified Version, together with at least five of the principal authors of the Document (all of its principal authors, if it has fewer than five), unless they release you from this requirement.

C. State on the Title page the name of the publisher of the Modified Version, as the

publisher. D. Preserve all the copyright notices of the Document. E. Add an appropriate copyright notice for your modifications adjacent to the other

copyright notices. F. Include, immediately after the copyright notices, a license notice giving the public

permission to use the Modified Version under the terms of this License, in the form shown in the Addendum below.

G. Preserve in that license notice the full lists of Invariant Sections and required Cover

Texts given in the Document’s license notice. H. Include an unaltered copy of this License. I. Preserve the section Entitled “History”, Preserve its Title, and add to it an item

stating at least the title, year, new authors, and publisher of the Modified Version as given on the Title Page. If there is no section Entitled “History” in the Document, create one stating the title, year, authors, and publisher of the Document as given on its Title Page, then add an item describing the Modified Version as stated in the previous sentence.

J. Preserve the network location, if any, given in the Document for public access to a Transparent copy of the Document, and likewise the network locations given in the Document for previous versions it was based on. These may be placed in the “History” section. You may omit a network location for a work that was published at least four years before the Document itself, or if the original publisher of the version it refers to gives permission.

K. For any section Entitled “Acknowledgements” or “Dedications”, Preserve the Title of the section, and preserve in the section all the substance and tone of each of the contributor acknowledgements and/or dedications given therein.

L. Preserve all the Invariant Sections of the Document, unaltered in their text and in their titles. Section numbers or the equivalent are not considered part of the section titles.

M. Delete any section Entitled “Endorsements”. Such a section may not be included

in the Modified Version. N. Do not retitle any existing section to be Entitled “Endorsements” or to conflict in

title with any Invariant Section. O. Preserve any Warranty Disclaimers.

If the Modified Version includes new front-matter sections or appendices that qualify as Secondary Sections and contain no material copied from the Document, you may at your option designate some or all of these sections as invariant. To do this, add their titles to the list of Invariant Sections in the Modified Version’s license notice. These titles must be distinct from any other section titles. You may add a section Entitled “Endorsements”, provided it contains nothing but endorsements of your Modified Version by various parties—for example, statements of peer review or that the text has been approved by an organization as the authoritative definition of a standard. You may add a passage of up to five words as a Front-Cover Text, and a passage of up to 25 words as a Back-Cover Text, to the end of the list of Cover Texts in the Modified Version. Only one passage of Front-Cover Text and one of Back-Cover Text may be added by (or through arrangements made by) any one entity. If the Document already includes a cover text for the same cover, previously added by you or by arrangement made by the same entity you are acting on behalf of, you may not add another; but you may replace the old one, on explicit permission from the previous publisher that added the old one. The author(s) and publisher(s) of the Document do not by this License give permission to use their names for publicity for or to assert or imply endorsement of any Modified Version.

5. COMBINING DOCUMENTS You may combine the Document with other documents released under this License, under the terms defined in section 4 above for modified versions, provided that you include in the combination all of the Invariant Sections of all of the original documents, unmodified, and list them all as Invariant Sections of your combined work in its license notice, and that you preserve all their Warranty Disclaimers. The combined work need only contain one copy of this License, and multiple identical Invariant Sections may be replaced with a single copy. If there are multiple Invariant Sections with the same name but different contents, make the title of each such section unique by adding at the end of it, in parentheses, the name of the original author or publisher of that section if known, or else a unique number. Make the same adjustment to the section titles in the list of Invariant Sections in the license notice of the combined work. In the combination, you must combine any sections Entitled “History” in the various original documents, forming one section Entitled “History”; likewise combine any sections Entitled “Acknowledgements”, and any sections Entitled “Dedications”. You must delete all sections Entitled “Endorsements.”

6. COLLECTIONS OF DOCUMENTS You may make a collection consisting of the Document and other documents released under this License, and replace the individual copies of this License in the various documents with a single copy that is included in the collection, provided that you follow the rules of this License for verbatim copying of each of the documents in all other respects. You may extract a single document from such a collection, and distribute it individually under this License, provided you insert a copy of this License into the extracted document, and follow this License in all other respects regarding verbatim copying of that document.

7. AGGREGATION WITH INDEPENDENT WORKS A compilation of the Document or its derivatives with other separate and independent documents or works, in or on a volume of a storage or distribution medium, is called an “aggregate” if the copyright resulting from the compilation is not used to limit the legal rights of the compilation’s users beyond what the individual works permit. When the Document is included in an aggregate, this License does not apply to the other works in the aggregate which are not themselves derivative works of the Document. If the Cover Text requirement of section 3 is applicable to these copies of the Document, then if the Document is less than one half of the entire aggregate, the Document’s Cover Texts may be placed on covers that bracket the Document within the aggregate, or the electronic equivalent of covers if the Document is in electronic form. Otherwise they must appear on printed covers that bracket the whole aggregate.

8. TRANSLATION Translation is considered a kind of modification, so you may distribute translations of the Document under the terms of section 4. Replacing Invariant Sections with translations requires special permission from their copyright holders, but you may include translations of some or all Invariant Sections in addition to the original versions of these Invariant Sections. You may include a translation of this License, and all the license notices in the Document, and any Warranty Disclaimers, provided that you also include the original English version of this License and the original versions of those notices and disclaimers. In case of a disagreement between the translation and the original version of this License or a notice or disclaimer, the original version will prevail. If a section in the Document is Entitled “Acknowledgements”, “Dedications”, or “History”, the requirement (section 4) to Preserve its Title (section 1) will typically require changing the actual title.

9. TERMINATION You may not copy, modify, sublicense, or distribute the Document except as expressly provided under this License. Any attempt otherwise to copy, modify, sublicense, or distribute it is void, and will automatically terminate your rights under this License. However, if you cease all violation of this License, then your license from a particular copyright holder is reinstated (a) provisionally, unless and until the copyright holder explicitly and finally terminates your license, and (b) permanently, if the copyright holder fails to notify you of the violation by some reasonable means prior to 60 days after the cessation. Moreover, your license from a particular copyright holder is reinstated permanently if the copyright holder notifies you of the violation by some reasonable means, this is the first time you have received notice of violation of this License (for any work) from that copyright holder, and you cure the violation prior to 30 days after your receipt of the notice. Termination of your rights under this section does not terminate the licenses of parties who have received copies or rights from you under this License. If your rights have been terminated and not permanently reinstated, receipt of a copy of some or all of the same material does not give you any rights to use it.

10. FUTURE REVISIONS OF THIS LICENSE The Free Software Foundation may publish new, revised versions of the GNU Free Documentation License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns.

See http://www.gnu.org/copyleft/.

Each version of the License is given a distinguishing version number. If the Document specifies that a particular numbered version of this License “or any later version” applies to it, you have the option of following the terms and conditions either of that specified version or of any later version that has been published (not as a draft) by the Free Software Foundation. If the Document does not specify a version number of this License, you may choose any version ever published (not as a draft) by the Free Software Foundation. If the Document specifies that a proxy can decide which future versions of this License can be used, that proxy’s public statement of acceptance of a version permanently authorizes you to choose that version for the Document.

11. RELICENSING “Massive Multiauthor Collaboration Site” (or “MMC Site”) means any World Wide Web server that publishes copyrightable works and also provides prominent facilities for anybody to edit those works. A public wiki that anybody can edit is an example of such a server. A “Massive Multiauthor Collaboration” (or “MMC”) contained in the site means any set of copyrightable works thus published on the MMC site. “CC-BY-SA” means the Creative Commons Attribution-Share Alike 3.0 license published by Creative Commons Corporation, a not-for-profit corporation with a principal place of business in San Francisco, California, as well as future copyleft versions of that license published by that same organization. “Incorporate” means to publish or republish a Document, in whole or in part, as part of another Document. An MMC is “eligible for relicensing” if it is licensed under this License, and if all works that were first published under this License somewhere other than this MMC, and subsequently incorporated in whole or in part into the MMC, (1) had no cover texts or invariant sections, and (2) were thus incorporated prior to November 1, 2008. The operator of an MMC Site may republish an MMC contained in the site under CC-BY-SA on the same site at any time before August 1, 2009, provided the MMC is eligible for relicensing.

ADDENDUM: How to use this License for your documents

To use this License in a document you have written, include a copy of the License in the document and put the following copyright and license notices just after the title page:

        Copyright (C)  year  your name.
        Permission is granted to copy, distribute and/or modify this document
        under the terms of the GNU Free Documentation License, Version 1.3
        or any later version published by the Free Software Foundation;
        with no Invariant Sections, no Front-Cover Texts, and no Back-Cover
        Texts.  A copy of the license is included in the section entitled ``GNU
        Free Documentation License''.

If you have Invariant Sections, Front-Cover Texts and Back-Cover Texts, replace the “with…Texts.” line with this:

          with the Invariant Sections being list their titles, with
          the Front-Cover Texts being list, and with the Back-Cover Texts
          being list.

If you have Invariant Sections without Cover Texts, or some other combination of the three, merge those two alternatives to suit the situation.

If your document contains nontrivial examples of program code, we recommend releasing these examples in parallel under your choice of free software license, such as the GNU General Public License, to permit their use in free software.

Key Index 51

Key Index

^

^[] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 ^] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 ^A . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 ^B, LEFT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 ^C - . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 ^C . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 ^C / . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 ^C = . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 ^C ^D A . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 ^C ^D c . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 ^C ^D C . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 ^C ^D f . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 ^C ^D h . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 ^C ^D i . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 ^C ^D I . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 ^C ^D l . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 ^C ^D L . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 ^C ^D p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 ^C ^D P . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 ^C ^D r . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 ^C ^D R A. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 ^C ^D R i. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 ^C ^D R u. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 ^C ^D s . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 ^C ^D S . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 ^C ^D u . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 ^C ~ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 ^C 2 a . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 ^C 2 b . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 ^C 2 m . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 ^C a . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 ^C B . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 ^C b - . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 ^C b . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 ^C b / . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 ^C b 2 a. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 ^C b 2 b. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 ^C b 2 m. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 ^C b a . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 ^C b C . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 ^C b C (for directories) . . . . . . . . . . . . . . . . . . . . 22 ^C b d . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 ^C b D . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 ^C b D (for directories) . . . . . . . . . . . . . . . . . . . . 22 ^C b e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 ^C b E . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 ^C b f Z. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 ^C b G . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 ^C b H . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 ^C b i . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 ^C b I . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 ^C b k . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 ^C b K . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 ^C b M . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 ^C b m D. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 ^C b m F. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 ^C b m G. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 ^C b m K. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 ^C b m M. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 ^C b m P. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 ^C b m R. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 ^C b N . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 ^C b o . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 ^C b O . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 ^C b R . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 ^C b R (for directories) . . . . . . . . . . . . . . . . . . . . 22 ^C b r G. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 ^C b r M. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 ^C b r O. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 ^C b S . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 ^C b T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 ^C b T (for directories) . . . . . . . . . . . . . . . . . . . . 22 ^C b V . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 ^C b x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 ^C b X . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 ^C b z . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 ^C b Z . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 ^C C (for directories). . . . . . . . . . . . . . . . . . . . . . 22 ^C c q . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 ^C c t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 ^C d . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 ^C D (for directories). . . . . . . . . . . . . . . . . . . . . . 22 ^C e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 ^C E . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 ^C ESC = . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 ^C f d . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 ^C f Z . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 ^C G . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 ^C H . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 ^C i . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 ^C I . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 ^C j . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 ^C J . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 ^C k . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 ^C K . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 ^C M . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 ^C m D . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 ^C m d, ^C m ^D . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 ^C m F . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 ^C m G . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 ^C m K . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 ^C m L . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 ^C m M . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 ^C m P . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 ^C m R . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 ^C m T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 ^C N . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 ^C n d . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 ^C n u . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 ^C o . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 ^C O . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 ^C p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 ^C P . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 ^C q . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 ^C R . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 ^C R (for directories). . . . . . . . . . . . . . . . . . . . . . 22 ^C r G . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 ^C r M . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 ^C r O . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 ^C s . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 ^C S . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 ^C t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 ^C T (for directories). . . . . . . . . . . . . . . . . . . . . . 22 ^C u . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 ^C U . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 ^C V . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 ^C W . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 ^C x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 ^C X . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 ^C z . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 ^C Z . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 ^E . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 ^F, RIGHT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 ^H, BKSPC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 ^K . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 ^L . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 ^O, ESC o . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 ^R, ^Xr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 ^S, ^Xs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 ^SPC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 ^U . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 ^W . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 ^X ^A . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 ^X ^F . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 ^X ^I . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 ^X ^L . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 ^X ^P . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 ^X ^R . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 ^X ^X . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 ^X 0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 ^X 1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 ^X 2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 ^X 4 a . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 ^X b m . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 ^X d, ^X ^D . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 ^X e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 ^X E . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 ^X g . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 ^X G . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 ^X h . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 ^X H . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 ^X I . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 ^X k . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 ^X N . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 ^X p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 ^X P . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 ^X q . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 ^X S . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 ^X T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 ^X v . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 ^X w . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 ^X W . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 ^X z . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 ^Y . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

D

DEL, ^D . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 DOWN, ^N . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

E

END, ESC > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 ESC % . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 ESC & . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 ESC + . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 ESC - . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 ESC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 ESC / . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 ESC \ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 ESC 5 (for directories) . . . . . . . . . . . . . . . . . . . . 22 ESC 6 (for directories) . . . . . . . . . . . . . . . . . . . . 22 ESC 8 (for directories) . . . . . . . . . . . . . . . . . . . . 22 ESC a c . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 ESC a o . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 ESC b . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 ESC BKSPC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 ESC c . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 ESC d . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 ESC e a . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 ESC e d . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 ESC e f . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 ESC e m . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 ESC e o . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 ESC e s . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 ESC e S . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 ESC ESC 1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 ESC ESC 2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 ESC ESC RET . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 ESC f . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 ESC g . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 ESC h . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 ESC i . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 ESC j . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 ESC k . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 ESC l . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 ESC m a . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 ESC m b . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 ESC m c . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 ESC m f . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 ESC m j . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 ESC m t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 ESC m z . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 ESC n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 ESC p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 ESC P a, ESC P l, ESC P u . . . . . . . . . . . . . . . . . . . . 29 ESC P b, ESC P c, ESC P e . . . . . . . . . . . . . . . . . . . . 29 ESC P T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 ESC P x, ESC P y . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 ESC r a . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 ESC r b . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 ESC r c . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 ESC r f . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 ESC r j . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 ESC r t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 ESC r z . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 ESC RET . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 ESC s d . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 ESC s e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 ESC S f . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 ESC s g i. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 ESC s g n. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 ESC s m . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 ESC S m . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 ESC s n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 ESC s o i. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 ESC s o n. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 ESC S q . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 ESC s s . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 ESC S s . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 ESC s u . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 ESC S u . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 ESC S v . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 ESC S w . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 ESC SPC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 ESC u . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 ESC w . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 ESC x r m. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

F

F1, ESC 1, ^X i . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 F10, ESC 0, ^X ^C, ^X c. . . . . . . . . . . . . . . . . . . . . . 30 F2, ESC 2, ^X a . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 F3, ESC 3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 F4, ESC 4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 F5 (for directories) . . . . . . . . . . . . . . . . . . . . . . . 22 F5, ESC 5, ^C C . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 F6 (for directories) . . . . . . . . . . . . . . . . . . . . . . . 22 F6, ESC 6, ^C T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 F7, ESC 7, ^X M . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 F8 (for directories) . . . . . . . . . . . . . . . . . . . . . . . 22 F8, ESC 8, ^C D . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 F9, ESC 9, ^X m . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

H

HOME, ESC < . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

I

INS, ^T, ^X \, ^\ . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

P

PGDOWN, ^V . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 PGUP, ESC v . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

T

TAB, ^I, ^X o . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

U

UP, ^P . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

Command Index 54

Command Index

A

adapt-current-directory . . . . . . . . . . . . . . . . . . . 23

Adapt-other-directory . . . . . . . . . . . . . . . . . . . . . 23 apropos. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 ASCII-MAIL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

B

B-ASCII-MAIL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 B-BINARY-MAIL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 B-BZIP2-COMPRESS . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 B-BZIP2-UNCOMPRESS . . . . . . . . . . . . . . . . . . . . . . . . . 15 B-CHGRP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 B-CHMOD. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 B-CHOWN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 B-COMPRESS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 B-COPY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

B-COPY (for directories) . . . . . . . . . . . . . . . . . . . 22

B-DELETE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

B-DELETE (for directories) . . . . . . . . . . . . . . . . . 22

B-DIRR-DIFF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 B-GENERIC-UNPACK . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 B-GENERIC-UNPACK-INTO . . . . . . . . . . . . . . . . . . . . . 18 B-LINK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 B-MAKE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 B-MIME-MAIL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 B-MIME-PACK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 B-MIME-UNPACK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 B-MOVE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

B-MOVE (for directories) . . . . . . . . . . . . . . . . . . . 22

B-MTOOLS-DELETE . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 B-MTOOLS-FORMAT . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 B-MTOOLS-GET . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 B-MTOOLS-MKDIR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 B-MTOOLS-PUT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 B-MTOOLS-RENAME . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 B-MTOOLS-RMDIR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 B-R-CHGRP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 B-R-CHMOD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 B-R-CHOWN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 B-RENAME. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

B-RENAME (for directories) . . . . . . . . . . . . . . . . . 22

B-RPM-INSTALL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 B-RPM-UNINSTALL . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 B-RPM-UPGRADE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 B-SPLIT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 B-SYMLINK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 B-TAR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 B-TAR-BZIP2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 B-TAR-COMPRESS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 B-UNCOMPRESS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 B-UUDECODE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 B-UUENCODE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 backward-char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 backward-delete-char. . . . . . . . . . . . . . . . . . . . . . . . 9 backward-kill-word . . . . . . . . . . . . . . . . . . . . . . . . . . 9 backward-word . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 beginning-of-line . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 beginning-of-panel . . . . . . . . . . . . . . . . . . . . . . . . . . 6 BF-UNCOMPRESS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 bin-packing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 BINARY-MAIL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 BZIP2-COMPRESS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 BZIP2-UNCOMPRESS . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

C

capitalize-word . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 change-directory . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 CHANGE-LOG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 CHGRP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 CHMOD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 CHOWN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 compare. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 COMPRESS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 copy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 copy (for directories) . . . . . . . . . . . . . . . . . . . . . 22

D

DATE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 DECRYPT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 delete . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

                                                delete (for directories) . . . . . . . . . . . . . . . . . . . 22

delete-char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

                                                delete-horizontal-space . . . . . . . . . . . . . . . . . . . . 9

DIFF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 DIRECTORY-USAGE . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 DIRR-DIFF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 DISK-FREE-SPACE . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 downcase-word . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 DPKG-CONFIGURE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 DPKG-CONTENTS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 DPKG-FIELD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 DPKG-HELP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 DPKG-INFO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 DPKG-INSTALL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 DPKG-LISTFILES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 DPKG-LISTPKGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 DPKG-PRINTAVAIL . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 DPKG-PURGE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 DPKG-RECORDAVAIL . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 DPKG-REMOVE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 DPKG-SEARCH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 DPKG-STATUS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 DPKG-UNPACK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

E

EDIT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 ENCRYPT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 end-of-line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 end-of-panel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 enlarge-other-panel . . . . . . . . . . . . . . . . . . . . . . . . . 4 enlarge-panel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 entry-to-input-line . . . . . . . . . . . . . . . . . . . . . . . . . 8 ENV . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

exchange-point-and-mark . . . . . . . . . . . . . . . . . . . 10

exit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30

F

F-UNCOMPRESS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 FAST-DIFF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 FILE-ACTION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 FILE-CREATE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 FILE-TYPE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 FIND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 FINGER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 FIRST-FLOPPY-DIR . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 forward-char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 forward-word . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

G

GENERIC-UNPACK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 GENERIC-UNPACK-INTO. . . . . . . . . . . . . . . . . . . . . . . . 18 GITPS, PS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 GREP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

H

HOME-DIR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

horizontal-scroll-left . . . . . . . . . . . . . . . . . . . . . 7
horizontal-scroll-right . . . . . . . . . . . . . . . . . . . . 7

HTML . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30

I

INCLUDE-DIR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 INFO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 isearch-backward . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 isearch-forward . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 ISPELL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

J

just-one-space . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

K

KILL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 kill-line. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 kill-region . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 kill-ring-save . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

kill-to-beginning-of-line . . . . . . . . . . . . . . . . . . 9

kill-to-end-of-line . . . . . . . . . . . . . . . . . . . . . . . . . 9 kill-word. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

L

LAST-BACKUP-DIFF . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 LINK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 LOCATE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 lock . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26

M

MAKE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 make-directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 MAN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 MIME-MAIL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 MOUNT-A. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 MOUNT-B. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 MOUNT-CDROM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 MOUNT-FLOPPY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 MOUNT-JAZ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 MOUNT-THESE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 MOUNT-ZIP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 MOUNTED-FILE-SYSTEMS. . . . . . . . . . . . . . . . . . . . . . . 28 move . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 move (for directories) . . . . . . . . . . . . . . . . . . . . . 22 MPACK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 MTOOLS-CHDIR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 MTOOLS-DELETE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 MTOOLS-DIR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 MTOOLS-FORMAT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 MTOOLS-GET . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 MTOOLS-MKDIR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 MTOOLS-PUT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 MTOOLS-RENAME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 MTOOLS-RMDIR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 MTOOLS-TYPE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 MULTIPLE-EDIT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 MULTIPLE-VIEW . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 MUNPACK. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

N

name-downcase . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 name-upcase . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 next-directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 next-history-element. . . . . . . . . . . . . . . . . . . . . . . 10 next-line. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

O

other-panel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

other-path-to-input-line . . . . . . . . . . . . . . . . . . . 8

P

panel-enable-abbrevsize . . . . . . . . . . . . . . . . . . . . 5

panel-enable-all . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

panel-enable-date-time . . . . . . . . . . . . . . . . . . . . . 5
panel-enable-full-name . . . . . . . . . . . . . . . . . . . . . 5

panel-enable-mode . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

panel-enable-next-mode . . . . . . . . . . . . . . . . . . . . . 5
panel-enable-owner-group . . . . . . . . . . . . . . . . . . . 5

panel-enable-size . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 panel-sort-by-date . . . . . . . . . . . . . . . . . . . . . . . . . . 6

panel-sort-by-extension . . . . . . . . . . . . . . . . . . . . 5
panel-sort-by-group-id . . . . . . . . . . . . . . . . . . . . . 6
panel-sort-by-group-name . . . . . . . . . . . . . . . . . . . 6

panel-sort-by-mode . . . . . . . . . . . . . . . . . . . . . . . . . . 6 panel-sort-by-name . . . . . . . . . . . . . . . . . . . . . . . . . . 5

panel-sort-by-owner-id . . . . . . . . . . . . . . . . . . . . . 6
panel-sort-by-owner-name . . . . . . . . . . . . . . . . . . . 6

panel-sort-by-size . . . . . . . . . . . . . . . . . . . . . . . . . . 6

panel-sort-next-method . . . . . . . . . . . . . . . . . . . . . 6

previous-directory . . . . . . . . . . . . . . . . . . . . . . . . . 24

previous-history-element . . . . . . . . . . . . . . . . . . 10

previous-line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 PRINT-JOB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 PRINT-JOB-LIST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 PSTREE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

Q

quick-compare-panels. . . . . . . . . . . . . . . . . . . . . . . 23 QUOTA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

R

R-CHGRP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 R-CHMOD. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 R-CHOWN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 R-DPKG-INSTALL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 R-DPKG-RECORDAVAIL . . . . . . . . . . . . . . . . . . . . . . . . . 19 R-DPKG-UNPACK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 READ-MAIL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 RECURSIVE-GREP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 refresh. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 RENAME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

RENAME (for directories) . . . . . . . . . . . . . . . . . . . 22
reset-directory-history . . . . . . . . . . . . . . . . . . . 24

ROOT-DIR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 RPM-INSTALL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 RPM-QUERY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 RPM-UNINSTALL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 RPM-UPGRADE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

S

scroll-down . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 scroll-up. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 SECOND-FLOPPY-DIR . . . . . . . . . . . . . . . . . . . . . . . . . . 24 select-entry . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 select-extension . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

select-files-matching-pattern . . . . . . . . . . . . . . 7

                                                selected-entries-to-input-line . . . . . . . . . . . . . 8

set-mark. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 set-scroll-step . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 SPLIT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 SUB-SHELL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 switch-panels . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 SYMLINK. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 SYNC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

T

TAR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 TAR-BZIP2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 TAR-COMPRESS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

                                                thoroughly-compare-panels . . . . . . . . . . . . . . . . . 23

tty-mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 TTY-RESET . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 two-panel-mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

U

UMOUNT-A. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 UMOUNT-B. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 UMOUNT-FLOPPY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 UMOUNT-JAZ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 UMOUNT-THESE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 UMOUNT-ZIP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 UNCOMPRESS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 unselect-extension . . . . . . . . . . . . . . . . . . . . . . . . . . 7

                                                unselect-files-matching-pattern . . . . . . . . . . . . 7

up-one-dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 upcase-word . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 USERS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 UUDECODE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 UUENCODE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

V

VIEW . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 VIRTUAL-MEMORY-STATUS . . . . . . . . . . . . . . . . . . . . . 28

W

WHEREIS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 WHICH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 WHO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 WIPE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

X

XHOST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

Y

yank . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

Variable Index 57

Variable Index

A

AnsiColors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35

C

CommandLineModeHelp. . . . . . . . . . . . . . . . . . . . . . . . 36 ConfirmOnExit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35

E

EDITOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

G

GNUIT_BROWSER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 GNUIT_EDITOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 GNUIT_RMAIL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 GNUIT_SHELL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 GNUIT_VMSTAT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 GroupDigits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35

H

HistoryFile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35

I

InfoDisplay . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35

L

LeadingDotMatch . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36

M

MaxUnscaledDigits . . . . . . . . . . . . . . . . . . . . . . . . . . 35

N

NormalModeHelp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36

S

SHELL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

                                                StartupFileDisplayMode . . . . . . . . . . . . . . . . . . . . 35

StartupFileSortMethod . . . . . . . . . . . . . . . . . . . . . 35 StartupScrollStep . . . . . . . . . . . . . . . . . . . . . . . . . . 35

T

TERM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 TypeSensitivity . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36

U

UseLastScreenChar . . . . . . . . . . . . . . . . . . . . . . . . . . 35

Concept Index 58

Concept Index

A

Adapt directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 Adding and removing hosts . . . . . . . . . . . . . . . . . . . 28 Archive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

B

Background directory copy . . . . . . . . . . . . . . . . . . . 22 Background directory delete . . . . . . . . . . . . . . . . . . 22 Background directory diff . . . . . . . . . . . . . . . . . . . . 23 Background directory move . . . . . . . . . . . . . . . . . . . 22 Background directory rename . . . . . . . . . . . . . . . . . 22 Background file copy . . . . . . . . . . . . . . . . . . . . . . . . . 11 Background file delete. . . . . . . . . . . . . . . . . . . . . . . . 11 Background file move . . . . . . . . . . . . . . . . . . . . . . . . 11 Background file rename . . . . . . . . . . . . . . . . . . . . . . 12 Background file split . . . . . . . . . . . . . . . . . . . . . . . . . 12 Background make. . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 Backward char. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 Backward delete char . . . . . . . . . . . . . . . . . . . . . . . . . 9 Backward kill word . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Backward word . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 bash . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 Beginning of line. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Beginning of panel . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Binary comparison . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 Binary files by mail . . . . . . . . . . . . . . . . . . . . . . . . . . 25 Browsing through the process list . . . . . . . . . . . . . 29 bzip2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

C

Capitalize word. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Change the current panel directory. . . . . . . . . . . . 23 Change the other panel directory . . . . . . . . . . . . . 23 Changing directory . . . . . . . . . . . . . . . . . . . . . . . . . . 23 Changing the current directory . . . . . . . . . . . . . . . 23 Changing the directory . . . . . . . . . . . . . . . . . . . . . . . 23 Changing the file’s group . . . . . . . . . . . . . . . . . . . . . 13 Changing the file’s mode . . . . . . . . . . . . . . . . . . . . . 13 Changing the file’s owner . . . . . . . . . . . . . . . . . . . . . 13 Changing the group in background . . . . . . . . . . . . 13 Changing the group recursively . . . . . . . . . . . . . . . 13 Changing the group recursively in background. . 13 Changing the inode’s group. . . . . . . . . . . . . . . . . . . 13 Changing the inode’s mode . . . . . . . . . . . . . . . . . . . 13 Changing the inode’s owner . . . . . . . . . . . . . . . . . . 13 Changing the mode in background . . . . . . . . . . . . 13 Changing the mode recursively . . . . . . . . . . . . . . . 13 Changing the mode recursively in background. . 13 Changing the owner in background . . . . . . . . . . . . 13 Changing the owner recursively . . . . . . . . . . . . . . . 13 Changing the owner recursively in background. . 13 chdir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 chgrp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 chmod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 chown . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 Comparing directories quickly . . . . . . . . . . . . . . . . 23 Comparing directories thoroughly . . . . . . . . . . . . . 23 Comparing panels quickly . . . . . . . . . . . . . . . . . . . . 23 Comparing panels thoroughly . . . . . . . . . . . . . . . . . 23 Compile. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 Compiling programs . . . . . . . . . . . . . . . . . . . . . . . . . 24 compress . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14, 15 Compressing directories . . . . . . . . . . . . . . . . . . . 14, 15 Compressing directories in background . . . . . 14, 15 Compressing files . . . . . . . . . . . . . . . . . . . . . . . . . 14, 15 Compressing files in background . . . . . . . . . . . 14, 15 Console lock. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 Convert to lower case . . . . . . . . . . . . . . . . . . . . . . . . 10 Convert to upper case . . . . . . . . . . . . . . . . . . . . . . . . 10 Copy entry name to input line. . . . . . . . . . . . . . . . . 8 Copying directories . . . . . . . . . . . . . . . . . . . . . . . . . . 22 Copying directories in background . . . . . . . . . . . . 22 Copying files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 Copying files in background . . . . . . . . . . . . . . . . . . 11 Create directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 Create file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 Creating a directory . . . . . . . . . . . . . . . . . . . . . . . . . 22 Creating bzip2 compressed tar archives . . . . . . . . 17 Creating bzip2 compressed tar archives in

background . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 Creating compressed tar archives. . . . . . . . . . . . . . 17 Creating compressed tar archives in background

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 Creating tar archives . . . . . . . . . . . . . . . . . . . . . . . . . 17 Creating tar archives in background . . . . . . . . . . . 17 csh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 Current disk quota. . . . . . . . . . . . . . . . . . . . . . . . . . . 28 Cursor backward. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 Cursor backward one word . . . . . . . . . . . . . . . . . . . . 9 Cursor down one entry . . . . . . . . . . . . . . . . . . . . . . . . 6 Cursor down one page. . . . . . . . . . . . . . . . . . . . . . . . . 6 Cursor end . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Cursor forward . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Cursor forward one word . . . . . . . . . . . . . . . . . . . . . . 9 Cursor home . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Cursor to BOL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Cursor to EOL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Cursor to the other panel. . . . . . . . . . . . . . . . . . . . . . 7 Cursor up one entry. . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Cursor up one page . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

D

date . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 Date and time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Date, time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Decoding files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Decoding files in background . . . . . . . . . . . . . . . . . 15 Decompressing directories . . . . . . . . . . . . . . . . . 14, 15 Decompressing directories in background … 14, 15 Decompressing directories/links . . . . . . . . . . . . . . . 15 Decompressing directories/links in background

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Decompressing files . . . . . . . . . . . . . . . . . . . . . . . 14, 15 Decompressing files in background . . . . . . . . . 14, 15 Decompressing files/links . . . . . . . . . . . . . . . . . . . . . 15 Decompressing files/links in background . . . . . . . 15 Decrypting the current file. . . . . . . . . . . . . . . . . . . . 16 Delete char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Delete horizontal space . . . . . . . . . . . . . . . . . . . . . . . . 9 Delete line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Delete spaces . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Delete word backward . . . . . . . . . . . . . . . . . . . . . . . . . 9 Delete word forward . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Deleting directories . . . . . . . . . . . . . . . . . . . . . . . . . . 22 Deleting directories in background . . . . . . . . . . . . 22 Deleting files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 Deleting files in background . . . . . . . . . . . . . . . . . . 11 diff . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 Differences between ascii files . . . . . . . . . . . . . . . . . 16 Differences between directories . . . . . . . . . . . . . . . . 23 Differences between file and backup . . . . . . . . . . . 16 Differences between two directories . . . . . . . . . . . . 23 Directory copy. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 Directory delete . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 Directory diff . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 Directory history reset . . . . . . . . . . . . . . . . . . . . . . . 24 Directory mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Directory move . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 Directory rename . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 Directory size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Directory usage. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 Disk free space . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 Downcase file names . . . . . . . . . . . . . . . . . . . . . . . . . 12 Downcase word . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 du . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

E

Editing a file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 Editing multiple files . . . . . . . . . . . . . . . . . . . . . . . . . 14 Editing the ChangeLog. . . . . . . . . . . . . . . . . . . . . . . 14 Emacs rmail. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 Enable all . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Enable next mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Encoding files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Encoding files in background . . . . . . . . . . . . . . . . . 15 Encrypting the current file . . . . . . . . . . . . . . . . . . . 16 End of line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 End of panel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Enlarge other panel . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Enlarge panel. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Entire screen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 env . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 Exchange point and mark . . . . . . . . . . . . . . . . . . . . 10 Exiting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 Extracting files from archives. . . . . . . . . . . . . . 17, 18

F

File action . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 File copy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 File create . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 File delete. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 File edit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 File mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 File move . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 File or directory mode . . . . . . . . . . . . . . . . . . . . . . . . 5 File or directory size . . . . . . . . . . . . . . . . . . . . . . . . . . 5 File rename . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 File search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 File size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 File split . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 File system status . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 File view . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 File wipe . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 find. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 finger . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 First entry . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Forward char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Forward word . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Free disk space . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 Full directory info. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Full directory name . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Full file info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Full file name. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Full file or directory name . . . . . . . . . . . . . . . . . . . . . 5 Full info. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Full name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

G

gitaction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 gitfm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 gitkeys . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 gitmount . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 gitps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 gitregrep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 gitrfgrep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 gitrgrep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 gitunpack . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 gitview. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 gitwipe . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 Go to .. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 Go to / . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 Go to the $HOME directory . . . . . . . . . . . . . . . . . . 24 Go to the /usr/include directory . . . . . . . . . . . . . . 24 Go to the ~ directory . . . . . . . . . . . . . . . . . . . . . . . . 24 Go to the first floppy mount point . . . . . . . . . . . . 24 Go to the home directory. . . . . . . . . . . . . . . . . . . . . 24 Go to the include directory . . . . . . . . . . . . . . . . . . . 24 Go to the other panel . . . . . . . . . . . . . . . . . . . . . . . . . 7 Go to the parent directory. . . . . . . . . . . . . . . . . . . . 24 Go to the root directory . . . . . . . . . . . . . . . . . . . . . . 24 Go to the second floppy mount point . . . . . . . . . . 24 Go up one directory. . . . . . . . . . . . . . . . . . . . . . . . . . 24 grep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 Group and owner . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Group, owner. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 gunzip. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14, 15 gzip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

H

Half screen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Hard links to files . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 Hot Keys . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

I

Incremental search backward . . . . . . . . . . . . . . . . . . 8 Incremental search forward . . . . . . . . . . . . . . . . . . . . 8 Isearch backward . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 Isearch forward . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 ispell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

J

Just one space . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

K

Kill line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Kill region . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Kill ring save . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Kill to beginning of line . . . . . . . . . . . . . . . . . . . . . . . 9 Kill to end of line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 kill word . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Killing processes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

L

Last backup diff . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 Last entry. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Linking files in background . . . . . . . . . . . . . . . . . . . 11 locate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 Locating files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 Locking the console . . . . . . . . . . . . . . . . . . . . . . . . . . 26 Logged in users . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 Looking at the environment . . . . . . . . . . . . . . . . . . 28 lpr. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

M

Major modes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 make . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 Make directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 Making a directory . . . . . . . . . . . . . . . . . . . . . . . . . . 22 Mark entry . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Mime encoded files by mail . . . . . . . . . . . . . . . . . . . 25 Minor modes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 mount . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 Mounted file systems list . . . . . . . . . . . . . . . . . . . . . 28 Mounting subdirectories . . . . . . . . . . . . . . . . . . . . . . 27 Mounting the cdrom . . . . . . . . . . . . . . . . . . . . . . . . . 27 Mounting the first floppy . . . . . . . . . . . . . . . . . 26, 27 Mounting the floppy . . . . . . . . . . . . . . . . . . . . . . 26, 27 Mounting the jaz drive . . . . . . . . . . . . . . . . . . . . . . . 27 Mounting the second floppy . . . . . . . . . . . . . . . . . . 26 Mounting the zip drive . . . . . . . . . . . . . . . . . . . . . . . 27 Mounting these file systems. . . . . . . . . . . . . . . . . . . 27 Moving directories . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 Moving directories in background . . . . . . . . . . . . . 22 Moving files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 Moving files in background . . . . . . . . . . . . . . . . . . . 11 mpack . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Multiple edit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 munpack . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

N

Next directory history entry . . . . . . . . . . . . . . . . . . 24 Next history element . . . . . . . . . . . . . . . . . . . . . . . . . 10 Next sort method . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 No panels mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

O

One panel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4, 5 Other panel path to input line . . . . . . . . . . . . . . . . . 8 Owner and group . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Owner, group. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

P

Pack files into bins . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 Previous directory history entry. . . . . . . . . . . . . . . 24 Previous history element . . . . . . . . . . . . . . . . . . . . . 10 printing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 Printing files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 Program search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

Q

Quota . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

R

Re-reading directories. . . . . . . . . . . . . . . . . . . . . . . . 26 Reading mail . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 Reading the html documentation. . . . . . . . . . . . . . 30 Reading the info documentation . . . . . . . . . . . . . . 29 Reading the manual pages . . . . . . . . . . . . . . . . . . . . 29 Recursive grep. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 Refreshing the screen. . . . . . . . . . . . . . . . . . . . . . . . . 26 regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 Renaming directories. . . . . . . . . . . . . . . . . . . . . . . . . 22 Renaming directories in background . . . . . . . . . . . 22 Renaming files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Renaming files in background . . . . . . . . . . . . . . . . . 12 Reset the directory history . . . . . . . . . . . . . . . . . . . 24 Resetting the terminal . . . . . . . . . . . . . . . . . . . . . . . 26

S

Save and delete region . . . . . . . . . . . . . . . . . . . . . . . . 9 Save into the kill ring . . . . . . . . . . . . . . . . . . . . . . . . . 9 Scroll left . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Scroll right . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Scroll step. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Searching binaries . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 Searching files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 Searching for key-bindings . . . . . . . . . . . . . . . . . . . . 30 Searching patterns in files . . . . . . . . . . . . . . . . . . . . 25 Searching programs . . . . . . . . . . . . . . . . . . . . . . . . . . 17 Searching regular expressions . . . . . . . . . . . . . . . . . 25 Select entry . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Selected entries to input line . . . . . . . . . . . . . . . . . . . 8 Selects files matching extension . . . . . . . . . . . . . . . . 7 Selects files matching pattern . . . . . . . . . . . . . . . . . . 7 Sending ascii mail . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 Sending ascii mail in background. . . . . . . . . . . . . . 25 Sending binary mail. . . . . . . . . . . . . . . . . . . . . . . . . . 25 Sending binary mail in background . . . . . . . . . . . . 25 Sending files by mail . . . . . . . . . . . . . . . . . . . . . . . . . 25 Sending mail . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 Sending mime mail . . . . . . . . . . . . . . . . . . . . . . . . . . 25 Sending mime mail in background. . . . . . . . . . . . . 25 Sending text by mail . . . . . . . . . . . . . . . . . . . . . . . . . 25 Set mark . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Set the mark . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 sh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 Shell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 Sorted by date. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Sorted by extension . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Sorted by group id . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Sorted by group name. . . . . . . . . . . . . . . . . . . . . . . . . 6 Sorted by mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Sorted by name. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Sorted by owner id. . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Sorted by owner name . . . . . . . . . . . . . . . . . . . . . . . . 6 Sorted by size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Sorting. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Sorting by date . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Sorting by extension . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Sorting by group id . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Sorting by group name . . . . . . . . . . . . . . . . . . . . . . . . 6 Sorting by mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Sorting by name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Sorting by owner id . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Sorting by owner name . . . . . . . . . . . . . . . . . . . . . . . . 6 Sorting by size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Spell checking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 Splitting files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Splitting files in background . . . . . . . . . . . . . . . . . . 12 Swapping . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 Switch panels. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Symbolic links to files . . . . . . . . . . . . . . . . . . . . . . . . 11 sync . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 Synchronizing file systems . . . . . . . . . . . . . . . . . . . . 29 System users . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

T

Tar archive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 tcsh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 Terminal reset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 Time and date. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Time, date . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Toggle flag . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Toggle panels . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 tty mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Two panel mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Two panels . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Type specific file action . . . . . . . . . . . . . . . . . . . . . . 22

U

umount . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 Uncompressing directories. . . . . . . . . . . . . . . . . 14, 15 Uncompressing directories in background … 14, 15 Uncompressing directories/links. . . . . . . . . . . . . . . 15 Uncompressing directories/links in background

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Uncompressing files. . . . . . . . . . . . . . . . . . . . . . . 14, 15 Uncompressing files in background . . . . . . . . . 14, 15 Uncompressing files/links. . . . . . . . . . . . . . . . . . . . . 15 Uncompressing files/links in background . . . . . . . 15 Unmark entry . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Unmounting subdirectories . . . . . . . . . . . . . . . . . . . 27 Unmounting the cdrom. . . . . . . . . . . . . . . . . . . . . . . 27 Unmounting the first floppy . . . . . . . . . . . . . . . . . . 27 Unmounting the floppy . . . . . . . . . . . . . . . . . . . . . . . 27 Unmounting the jaz drive . . . . . . . . . . . . . . . . . . . . 27 Unmounting the second floppy . . . . . . . . . . . . . . . . 27 Unmounting the zip drive . . . . . . . . . . . . . . . . . . . . 27 Unmounting these file systems . . . . . . . . . . . . . . . . 27 Unselects files matching extension . . . . . . . . . . . . . . 7 Unselects files matching pattern. . . . . . . . . . . . . . . . 7 Upcase file names . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Upcase word . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 uudecode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 uuencode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

V

Viewing files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 Viewing multiple files . . . . . . . . . . . . . . . . . . . . . . . . 14 Viewing processes. . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 Viewing the process tree . . . . . . . . . . . . . . . . . . . . . 29 Virtual memory status . . . . . . . . . . . . . . . . . . . . . . . 28

W

whereis . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 which . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 who . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 Wiping files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 Wrapped incremental search . . . . . . . . . . . . . . . . . . . 8 Wrapped isearch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 Wrapped search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

X

xhost . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

Y

Yanking, reinserting. . . . . . . . . . . . . . . . . . . . . . . . . . 10

Short Contents

1 Introduction. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 Distributing GNU Interactive Tools . . . . . . . . . . . . . . . . . . . 2 3 Using GNU Interactive Tools . . . . . . . . . . . . . . . . . . . . . . . 3 4 Customizing GNU Interactive Tools . . . . . . . . . . . . . . . . . . 33 5 GNU Interactive Tools limitations . . . . . . . . . . . . . . . . . . . 41 6 GNU Interactive Tools bugs . . . . . . . . . . . . . . . . . . . . . . . 42 A Copying This Manual . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 Key Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 Command Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 Variable Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 Concept Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58

Table of Contents

1 Introduction. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

2 Distributing GNU Interactive Tools . . . . . . . . . 2

3 Using GNU Interactive Tools . . . . . . . . . . . . . . . 3

3.1 The GIT file system browser . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3.1.1 Key binding conventions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3.1.2 Command line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 3.1.3 Panel modes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 3.1.4 Sorting methods . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 3.1.5 Moving the cursor in the panel. . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.1.6 Selecting files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 3.1.7 Incremental searching files in a panel . . . . . . . . . . . . . . . . . . . . . 8 3.1.8 Using the input line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

3.1.8.1 Inserting Text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 3.1.8.2 Moving Point . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 3.1.8.3 Deleting and killing text . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 3.1.8.4 Case conversion of words. . . . . . . . . . . . . . . . . . . . . . . . . . . 10 3.1.8.5 Reusing recent input line arguments . . . . . . . . . . . . . . . . 10 3.1.8.6 Commands to set the mark . . . . . . . . . . . . . . . . . . . . . . . . 10 3.1.8.7 Reinserting recently killed text . . . . . . . . . . . . . . . . . . . . . 10 3.1.8.8 Selecting files matching patterns. . . . . . . . . . . . . . . . . . . . 10

3.1.9 File operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 3.1.9.1 Copying Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 3.1.9.2 Moving Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 3.1.9.3 Creating Files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 3.1.9.4 Deleting Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 3.1.9.5 Linking Files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 3.1.9.6 Renaming Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 3.1.9.7 Splitting files into smaller parts . . . . . . . . . . . . . . . . . . . . 12 3.1.9.8 Packing files into the minimum number of bins . . . . . . 12 3.1.9.9 Changing a file’s mode, owner and group. . . . . . . . . . . . 13 3.1.9.10 Editing Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 3.1.9.11 Viewing Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 3.1.9.12 Compressing Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 3.1.9.13 Encoding Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 3.1.9.14 Encrypting Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 3.1.9.15 Comparing Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 3.1.9.16 Spell Checking Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 3.1.9.17 Printing Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 3.1.9.18 Wiping Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 3.1.9.19 Searching Files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 3.1.9.20 Managing tar based archive files . . . . . . . . . . . . . . . . . . . 17 3.1.9.21 Working with DPKG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 3.1.9.22 Installing and Uninstalling RPM packages. . . . . . . . . . 19 3.1.9.23 File Types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 3.1.9.24 Accessing MSDOS Files . . . . . . . . . . . . . . . . . . . . . . . . . . 20 3.1.9.25 A different action for each file type . . . . . . . . . . . . . . . . 21

3.1.10 Directory operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 3.1.10.1 Creating directories . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 3.1.10.2 Copying directories . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 3.1.10.3 Deleting directories. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 3.1.10.4 Moving directories . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 3.1.10.5 Renaming directories . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 3.1.10.6 Comparing Directories. . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 3.1.10.7 Summarize directory usage. . . . . . . . . . . . . . . . . . . . . . . . 23 3.1.10.8 Changing directories. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 3.1.10.9 Directory History . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 3.1.10.10 Hot Keys . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

3.1.11 Compiling programs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 3.1.12 Sending/receiving ascii/binary mail . . . . . . . . . . . . . . . . . . . . 25 3.1.13 Starting a sub-shell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 3.1.14 Using grep and recursive grep. . . . . . . . . . . . . . . . . . . . . . . . . . 25 3.1.15 Locking your console . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 3.1.16 Refreshing the screen contents . . . . . . . . . . . . . . . . . . . . . . . . . 26 3.1.17 Resetting your terminal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 3.1.18 Mounting/unmounting file systems . . . . . . . . . . . . . . . . . . . . . 26 3.1.19 Getting some useful system information . . . . . . . . . . . . . . . . 28 3.1.20 How to look at the environment variables . . . . . . . . . . . . . . . 28 3.1.21 Viewing/killing processes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 3.1.22 Synchronizing the file systems . . . . . . . . . . . . . . . . . . . . . . . . . 29 3.1.23 Reading the documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 3.1.24 Exiting GNU Interactive Tools. . . . . . . . . . . . . . . . . . . . . . . . . 30

3.2 The GNUIT process viewer/killer. . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 3.3 The GNUIT ASCII/HEX file viewer . . . . . . . . . . . . . . . . . . . . . . . . . 30 3.4 The GNUIT key sequences display utility . . . . . . . . . . . . . . . . . . . . 31 3.5 The GNUIT wipe file utility. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 3.6 The GNUIT mount utility . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 3.7 The GNUIT per file type action script . . . . . . . . . . . . . . . . . . . . . . . 31 3.8 Unified archive unpacking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 3.9 The GIT recursive grep script . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32

4 Customizing GNU Interactive Tools . . . . . . . 33

4.1 Environment Variables. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 4.2 Configuration Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

4.2.1 Writing key sequences . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 4.2.2 The global setup section. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 4.2.3 gitfm Sections . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35

4.2.3.1 gitfm Setup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 4.2.3.2 Using gitfm on color displays . . . . . . . . . . . . . . . . . . . . . . . 36 4.2.3.3 Using gitfm on monochrome displays . . . . . . . . . . . . . . . 36 4.2.3.4 Defining keys . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 4.2.3.5 The key-sequence field . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 4.2.3.6 The command-name field . . . . . . . . . . . . . . . . . . . . . . . . . . 37 4.2.3.7 The formatted-command field . . . . . . . . . . . . . . . . . . . . . . 37 4.2.3.8 The %s format specifier. . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 4.2.3.9 The %f format specifier . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 4.2.3.10 The %d format specifier . . . . . . . . . . . . . . . . . . . . . . . . . . 37 4.2.3.11 The %l format specifier . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 4.2.3.12 The %t format specifier. . . . . . . . . . . . . . . . . . . . . . . . . . . 37 4.2.3.13 The %z format specifier . . . . . . . . . . . . . . . . . . . . . . . . . . 37 4.2.3.14 The %a format specifier . . . . . . . . . . . . . . . . . . . . . . . . . . 37 4.2.3.15 The %m format specifier. . . . . . . . . . . . . . . . . . . . . . . . . . 38 4.2.3.16 The %g format specifier . . . . . . . . . . . . . . . . . . . . . . . . . . 38 4.2.3.17 The %o format specifier . . . . . . . . . . . . . . . . . . . . . . . . . . 38 4.2.3.18 The %p format specifier . . . . . . . . . . . . . . . . . . . . . . . . . . 38 4.2.3.19 The %b format specifier . . . . . . . . . . . . . . . . . . . . . . . . . . 38 4.2.3.20 The %i format specifier . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 4.2.3.21 The %? format specifier . . . . . . . . . . . . . . . . . . . . . . . . . . 38 4.2.3.22 The new-dir field . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 4.2.3.23 The save-screen field . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 4.2.3.24 The pause field . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 4.2.3.25 The hide field . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39

4.2.4 Setting up colors for different file types . . . . . . . . . . . . . . . . . . 39 4.2.5 gitps Sections . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39

4.2.5.1 gitps Setup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 4.2.5.2 Using gitps on color displays . . . . . . . . . . . . . . . . . . . . . . . 39 4.2.5.3 Using gitps on monochrome displays . . . . . . . . . . . . . . . . 39 4.2.5.4 Defining keys . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39

4.2.6 gitview Sections . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 4.2.6.1 gitview Setup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 4.2.6.2 Using gitview on color displays . . . . . . . . . . . . . . . . . . . . . 40 4.2.6.3 Using gitview on monochrome displays. . . . . . . . . . . . . . 40 4.2.6.4 Defining keys . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40

5 GNU Interactive Tools limitations . . . . . . . . . 41

6 GNU Interactive Tools bugs . . . . . . . . . . . . . . . 42 Appendix A Copying This Manual . . . . . . . . . . 43 Key Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 Command Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 Variable Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 Concept Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58